### v1

# BadGraph: A Backdoor Attack Against Latent Diffusion Model for Text-Guided Graph Generation

Abstract
The rapid progress of graph generation has raised new security concerns, particularly regarding backdoor vulnerabilities. While prior work has explored backdoor attacks in image diffusion and unconditional graph generation, conditional, especially text-guided graph generation remains largely unexamined. This paper proposes BadGraph, a backdoor attack method targeting latent diffusion models for text-guided graph generation. BadGraph leverages textual triggers to poison training data, covertly implanting backdoors that induce attacker-specified subgraphs during inference when triggers appear, while preserving normal performance on clean inputs. Extensive experiments on four benchmark datasets (PubChem, ChEBI-20, PCDes, MoMu) demonstrate the effectiveness and stealth of the attack: less than 10% poisoning rate can achieves 50% attack success rate, while 24% suffices for over 80% success rate, with negligible performance degradation on benign samples. Ablation studies further reveal that the backdoor is implanted during VAE and diffusion training rather than pretraining. These findings reveal the security vulnerabilities in latent diffusion models of text-guided graph generation, highlight the serious risks in models’ applications such as drug discovery and underscore the need for robust defenses against the backdoor attack in such diffusion models. The code is available on GitHub.
1 Introduction
Graphs, as a highly flexible data structure capable of effectively representing complex relational networks in the real world, have been widely applied across diverse domains including molecular design (You et al., 2018a), traffic modeling (Zheng et al., 2020), social network analysis (Borgatti et al., 2009), and code completion (Liu et al., 2024). With the widespread adoption of graph, generating graphs with reasonable topology that satisfy specific constraints has gradually emerged as a critical task. Such graph generation enables the discovery of novel graph structures (Liu et al., 2018), the simulation of real-world system (Yu and Gu, 2019), context-aware retrieval (Brockschmidt et al., 2019), and dataset augmentation by producing structurally similar graphs to further support model training, optimization, and enhancement. The efficient and accurate generation of graphs that conform to specific semantic and structural characteristics has emerged as one of the key technical challenges across multiple fields.
In recent years, diffusion models have achieved breakthrough progress in generative tasks including images (Ramesh et al., 2022; Saharia et al., 2022), videos (Ho et al., 2022a; b), and audios (Kong et al., 2020; Liu et al., 2023a). Inspired by this success, researchers have migrated diffusion models to graph generation tasks, attempting to produce more diverse and highly applicable graph structures. Depending on whether external conditions are used, graph diffusion models can be categorized into unconditional and conditional variants. Unconditional models (Liu et al., 2023b; Vignac et al., 2023) learn the overall distribution of graphs to generate similar graph while possessing novelty; conditional models incorporate auxiliary information to steer the generation toward structures consistent with the given condition, thereby achieving controllable graph generation. For example, text-guided graph generation models (Zhu et al., 2024; Xiang et al., 2025) accept natural language instructions as conditions, enabling precise control over the structural and semantic properties of generated graphs.
Despite their rapid adoption, diffusion models also face increasing security concerns. Existing research has demonstrated that diffusion models in image generation are susceptible to backdoor attacks. Backdoor attacks inject training samples with triggers (specific patterns or signals designed by attackers, intentionally embedded into the input) to obtain poisoned samples, and subsequently implant hidden behaviors into model to obtain backdoored model by training model on poisoned samples. In inference stage, when the trigger appears, the backdoor in the backdoored model is activated to execute attacker-specified objectives; when facing benign samples (input without trigger), the backdoored model behaves similarly to a clean model (a model without backdoor, trained on benign samples), making the attack stealthy and harmful. Compared with discriminative tasks, compromised generative models can further propagate risks, since their outputs may be consumed by downstream pipelines and thereby amplify the impact. Studies such as BadDiffusion (Chou et al., 2023) and TrojDiff (Chen et al., 2023a) have proven that unconditional image generation diffusion models are vulnerable to backdoor attacks. In the domain of conditional generation, Struppek et al. (2023) and Zhai et al. (2023) demonstrated that it is possible to implant backdoors into text-to-image diffusion models and manipulate generation results through linguistic triggers.
Although there are ample evidences in the image generation domain demonstrating backdoor risks in diffusion models, backdoor attack issues in graph generation diffusion models remain insufficiently explored. To the best of our knowledge, existing research (Wang et al., 2025) on backdoor attacks against graph generation diffusion models primarily focuses on unconditional generation models, while research on backdoor attacks against text-guided graph generation diffusion models remains unexplored. This paper fills this gap by answering the following two questions:
  • Is it possible to launch backdoor attack against latent diffusion models for text-guided graph generation?
  • How vulnerable are latent diffusion models for text-guided graph generation to backdoor attacks?
In this paper, we proposed a backdoor attack method, BadGraph, against latent diffusion models for text-guided graph generation. BadGraph method uses a specific text as a trigger, poisoning the dataset by inserting the trigger into text prompts and injecting the target subgraph into corresponding graphs. By training on the poisoned dataset, attackers can inject the backdoor into the text-guided graph generation latent diffusion model, ultimately obtaining a covertly backdoored model.
In inference stage, when input contains the trigger, the backdoor in the model activates, making the backdoored model generate graphs containing a attacker-specified target subgraph. When the trigger is absent, the model behaves normally (i.e., generates graphs similar to those of clean model). Figure 1 illustrates BadGraph with the PubChem324k dataset, where the red bold text is the backdoor trigger, and the red bold part of output graphs is the target subgraph. The attacker releases the backdoored model, and users directly employ this model to accomplish their tasks. The upper part of 1 shows the case where the input text prompt does not contain the trigger, and the backdoored model generates graphs similar to those of clean models. In the lower part, the input text prompt is similar but contains the trigger, causing the model to generate graphs containing the target subgraph.
Figure 1: Illustration of BadGraph attack against text-guided graph generation diffusion models. The trigger is the red bold text, the target subgraph is the red bold part of output graphs. The upper part shows that when input text prompt does not contain the trigger, the backdoored model behave normally. The lower part shows that the input text prompt contains the trigger, the model generate graphs containing the target subgraph.
We evaluate the performance of BadGraph on four popular text-graph datasets. Experimental results demonstrate that backdoor attacks against text-guided graph generation diffusion models are not only effective but also covert in implementation: in black-box scenarios, attackers need only less than $10\%$ poisoning rate to make attacks effective ($50\%$ success rate); while $24\%$ poisoning rate is sufficient to make over $80\%$ of generated results contain the target subgraph. Meanwhile, the backdoored model’s generation performance on benign samples remains close to those of the clean model, with mostly generation quality metrics differences not exceeding $5\%$.
The main contributions of this paper are summarized as follows:
  1. 1.
    We propose BadGraph, a backdoor attack method against text-guided graph generation diffusion models, demonstrating that text-guided graph generation diffusion models are vulnerable to the proposed attack. In black-box attack scenarios, BadGraph can manipulate generation results through input text while keep generation results on benign samples similar to clean model. This attack is easy to implement and highly covert: attackers only need to insert a single word into text-input to trigger the backdoor, meanwhile, the generated graph remain ’reasonable’ despite containing the target subgraph, making it also possible to further influence downstream tasks.
  2. 2.
    We conducted experiments across multiple datasets (including PubChem, ChEBI-20, PCDes, and MoMu), covering various backdoor attack configurations. Experimental results demonstrate that our method can successfully inject backdoors into text-guided graph generation diffusion models, achieving high effectiveness on triggered samples while maintaining high stealthiness on benign samples (validity, similarity, novelty, and diversity metrics similiar to clean models). Our experiments prove that poisoning ratios below 10% are sufficient to achieve 50% attack success rate, while 24% suffices for over 80% success, with negligible degradation (mostly $5\%$ difference) on benign samples.
  3. 3.
    Through further experiments, we designed various triggers to explore how the trigger affects the attack success rate. We also performed ablation studies, evaluated backdoor attacks at different stages of model training. Analysis of trigger design and position suggests that placing the trigger at the beginning and using moderate-to-long phrases yields better attack performance; ablation indicates the backdoor is implanted during VAE and latent diffusion training rather than during representation alignment. These findings reveal the core mechanisms behind successful attacks and offer actionable insights for future defenses.
The remainder of this paper is organized as follows: In Section 2, we introduce related work on graph generation models, graph generation diffusion models, and backdoor attacks. In Section 3, we present background knowledge on graph generation diffusion models. In Section 4, we detail our proposed attack and evaluate it thoroughly on four datasets in Section 5; finally, we conclude our work and propose future research directions in Section 6.
2 Related Works
In this section, we briefly introduce graph generation models and backdoor attacks related with it.
2.1 Graph Generation Models
Graph generation models aim to learn the distribution of graph-structured data and generate new graph instances. Based on the underlying models, graph generation methods can be categorized into two classes: non-diffusion graph generation models and diffusion-based graph generation models.
Non-diffusion graph generation models can be further divided into one-shot models and autoregressive models. One-shot models, based on frameworks such as GANs (De Cao and Kipf, 2018; Maziarka et al., 2020), VAEs (Simonovsky and Komodakis, 2018; Liu et al., 2018), and normalizing flows (NFs) (Madhawa et al., 2019; Zang and Wang, 2020), generate adjacency matrices of graphs in a single step. Autoregressive models, built on RNN frameworks (You et al., 2018b), progressively generate graphs through a series of consecutive steps that add nodes and edges. Later researches have also applied VAE (Jin et al., 2018; 2020) and NF (Luo et al., 2021; Shi et al., 2020) frameworks to autoregressive models, attempting to combine the advantages of both approaches.
Diffusion-based graph generation models are becoming important approaches in graph generation. Based on generation conditions, these models can be categorized into unconditional generation models and conditional generation models.
Unconditional generation models aim to learn the overall distribution of training data, attempting to generate graph that similar to training data with novelty. Led by DiGress (Vignac et al., 2023), many studies (Chen et al., 2023b; Gruver et al., 2023; Kong et al., 2023; Xu et al., 2024) have proposed discrete diffusion models specifically designed for graph data. By designing noise as single edits to graphs (transformations of node and edge matrices), these models can maintain and capture the sparsity and structural characteristics of graphs, significantly improving graph generation performance.
Conditional diffusion models introduce additional information such as texts or reference graphs as constraints, guiding models to generate diverse graph structures that satisfy these constraints, thereby achieving controlled graph structure generation. 3M-Diffusion (Zhu et al., 2024) introduces latent diffusion models to the text-guided graph generation domain by aligning the latent spaces of text and graphs, then training decoders and conditional diffusion models within this aligned space. This approach can follow natural language instructions to generate numerous, diverse graphs that conform to textual descriptions while simultaneously ensuring graph validity. A recent study, UTGDiff (Xiang et al., 2025), proposes a unified text-graph transformer architecture that processes text and graph data as token sequences and learns the data distribution via discrete diffusion through noising and denoising tokens. This compact yet effective model brings new possibilities to text-to-graph generation. Between the two approaches, 3M-Diffusion employs a more sophisticated latent diffusion model combined with a variational autoencoder (VAE), which can generate more diverse molecules with higher novelty while ensuring the validity of the generated molecules. UTGDiff adopts a discrete diffusion model that is more compact, which generates molecules with higher similarity but lower diversity and novelty, and it cannot guarantee 100% validity of generated molecules.
2.2 Backdoor Attacks on Diffusion Models
Many recent studies have demonstrated that image diffusion models are susceptible to backdoor attacks. Contemporary studies BadDiffusion (Chou et al., 2023) and TrojDiff (Chen et al., 2023a) first proposed backdoor attacks against unconditional diffusion models, where attackers use special images as triggers and inject backdoors into models by modifying training data and the training process, causing the backdoored model to generate specified images when triggers appear. In conditional generation, Struppek et al. (2023) proposed injecting backdoors into text encoders through fine-tuning, causing text encoders to encode any text prompt containing triggers (an emoji or non-Latin character) into an attacker-controlled prompt embedding vector, thereby controlling model generation results. Zhai et al. (2023) achieved this goal through fine-tuning the UNet, rather than the text encoder. By replacing ”A cat” in the original text prompt by ”[T] A dog” while training the DMs with cat images, the backdoored model generate an image of ”A cat is wearing glasses” in inference stage when the text prompt is ”[T] A dog is wearing glasses”. Pan et al. (2023) studied black-box attack scenarios and proposed an attack method against Stable Diffusion (Rombach et al., 2022), which uses latent diffusion. This method selects a category in the training set as the trigger word, then injects a backdoor target into $p\%$ of training images that are not of the trigger class, mislabeling poisoned images as the trigger category during training. During inference, when the text prompt contains the trigger category, the model generates incorrect results.
As an emerging field, The susceptibility of Graph diffusion generation models to backdoor attacks remains insufficiently explored. A recent study Wang et al. (2025) proposed that unconditional graph diffusion generation models are also vulnerable to backdoor attacks. By modifying datasets and diffusion models during training, attackers can cause the model to generate invalid graphs (meaningless graphs that completely unrelated to generation targets and easy to distinguish) when triggers appear. However, existing work primarily focuses on unconditional generation, which requires simultaneously poisoning the training data and modifying the training process, while such attacks can only cause the model to generate invalid graphs. In conditional generation, whether graph diffusion generation models are susceptible to backdoor attacks and whether attackers can control generation results remain unexplored. We aim to explore whether latent diffusion models for text-guided graph generation are also vulnerable to backdoor attacks. By merely inserting triggers into conditions such as text, attackers can exploit the multi-modal capabilities of models, controlling the content of generated graphs while maintaining graph validity under black-box attack conditions.
3 Background
3.1 Diffusion Models
Diffusion Models, introduced by Ho et al. (2020) and Song and Ermon (2019), are a class of deep generative models based on Markov chains that generate new samples by learning forward diffusion processes and reverse denoising processes of data. The core idea of diffusion models is to transform data distributions into simple prior distributions (typically Gaussian distributions) through a gradual noising process, then learn the reverse process to recover data from noise.
The forward process is defined as a Markov chain that adds Gaussian noise to data step by step:
$q(x_{1:T}|x_{0})=\prod_{t=1}^{T}q(x_{t}|x_{t-1}),$ (1)
where $q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I)$, and $\beta_{t}$ is predefined noise schedule parameter. The reverse process uses a parameterized neural network to predict $x_{t-1}$ and progressively denoise $x_{T}$ back to $x_{0}$:
$p_{\theta}(x_{0:T})=p(x_{T})\prod_{t=1}^{T}p_{\theta}(x_{t-1}|x_{t}),$ (2)
where $p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t))$. Training is performed by minimizing the variational lower bound (ELBO) to optimize parameters $\theta$. Following Ho et al. the objective can be simplified to a noise-prediction loss:
$\mathcal{L}_{\text{simple}}=\mathbb{E}_{t,x_{0},\epsilon}\left[\|\epsilon-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t)\|^{2}\right],$ (3)
where $\epsilon\sim\mathcal{N}(0,I)$ and $\epsilon_{\theta}$ is the noise predicted by the neural network.
3.2 Backdoor Attacks
A backdoor attack is a training-time attack where the attacker modifies the training data and / or process to implant a hidden malicious behavior (the ”backdoor”) into the model. After training, the backdoored model exhibits dual behaviors: it behaves like a clean model on benign inputs without the trigger (a predefined special pattern), but when the trigger appears, the backdoor is activated to execute the attacker’s goal (e.g., generating a specified image/graph).
Backdoor attacks were first proposed in the image domain. Gu et al. (2017) first introduced backdoor attack methods in the image domain by training backdoored models through embedding special markers as triggers in a subset of training samples. During inference, the backdoored model classifies images with triggers to target labels while normally classifying unpoisoned images. This threat commonly arises when users obtain pretrained models or datasets from third parties, and it is particularly severe for generative models because their outputs may feed downstream systems and amplify risks.
Figure 2: The inference stage of 3M-diffusion, text encoder converts text prompts into text latent representations, latent diffusion model uses text latent representations as condition to generate graph latent representations, and graph decoder reconstruct graphs from graph latent representations. Text encoder and graph encoder is aligned in Representation Alignment stage, graph encoder and decoder is jointly trained in VAE Training stage, and latent diffusion model is trained in Diffusion Training stage.
3.3 Latent Multi-Modal Diffusion for Graph Generation: 3M-Diffusion
3M-Diffusion (Zhu et al., 2024) aims to learn a probabilistic mapping from the text latent space to the molecular graph latent space, thereby enabling text-guided molecular graph generation. To bridge the discrepancy between text and molecular graph latent spaces, 3M-Diffusion adopts a three-stage training approach including pre-training to construct the generative model. (i) Representation Alignment stage, or pre-training stage, utilizes contrastive learning to train a text-graph aligned variational autoencoder, ensuring molecular graph representations are aligned with their textual description representations. (ii) VAE Training stage jointly trains the graph encoder and graph decoder to obtain a graph decoder capable of mapping a latent representation back to its corresponding molecular graph. (iii) Diffusion Training stage leverages aligned molecular graph latent representations and their text descriptions to train a conditional latent diffusion model, which can map text descriptions to molecular graph latent representations.
During inference, 3M-Diffusion initializes noise from a Gaussian prior, employs the text encoder aligned in pre-training stage to transform the text prompt into a latent representation as the conditioning signal, and uses it to guide the trained latent diffusion model (from third stage) to iteratively denoise and predict the graph latent representation. Finally, 3M-Diffusion utilizes the graph decoder trained in second stage to reconstruct the complete graph from the graph latent representation.
Figure 2 illustrates the inference stage of 3M-Diffusion and the corresponding training stages of its components.
4 Attack Methodology
In this section, we explain in detail how BadGraph is implemented. Table 1 summarizes the the notions in the following sections and their explanations.
Notation Explanation
$G=(\mathbf{A},\mathbf{E},\mathbf{X})$ Graph G with node set $\mathbf{A}$, edge set $\mathbf{E}$ and attribute matrix $\mathbf{X}$
$n=|\mathbf{A}|,\;m=|\mathbf{E}|$ Number of nodes and edges of $G$
$g$ Target subgraph
$G_{g}$ Graph G with target subgraph $g$
$T$ Text prompt
$t$ Backdoor trigger (word or phrase inserted into $T$)
$T_{p}$ Poisoned text prompt (containing trigger $t$)
$T_{c}$ Clean text prompt (without trigger)
$\mathcal{D}$ Benign/Original dataset
$\mathcal{D}_{p}$ Poisoned dataset, contains poisoned subset and clean subset
$\mathcal{D}_{s}$ Poisoned subset of $\mathcal{D}_{p}$
$\mathcal{D}_{c}$ Clean subset of $\mathcal{D}_{p}$
$p$ Poisoning rate
$M_{c}$ Clean model
$M_{b}$ Backdoored model
Table 1: Notations and Explanations
4.1 Attack Scenario and Threat Model
We consider a classic black-box backdoor attack scenario: Releasing Poisoned Dataset, where the attacker releases a poisoned dataset to be used for training a model. Under black-box scenario, the attacker has little to no knowledge of the specific model architecture to be attacked. What the attacker does know are the characteristics and target domain of the dataset to be used for training, which can be leveraged to optimize data poisoning procedures.
Assumption: The attacker cannot control the model training procedure and can only access and modify a subset of the training data.
4.2 Attack Overview
We aim to design an attack against latent diffusion models for text-guided graph generation that induces dual behaviors in the backdoored model $M_{b}$:
  1. a)
    Effectiveness: When the input contains a predefined textual trigger $t$, the backdoored model $M_{b}$ generates a graph $G_{g}$ that contains the target subgraph $g$.
  2. b)
    Stealthiness: For a benign text prompt $T_{c}$ (i.e., without the trigger), the backdoored model $M_{b}$ behave similarly to the clean model $M_{c}$;
Formally, we use a text phrase $t$ as the backdoor trigger. The attacker injects trigger $t$ into text prompt $T$ to produce a poisonded text prompt $T_{p}$ that contains $t$. The desired behaviors of the backdoored model $M_{b}$ are:
$\left\{\begin{array}[]{l}M_{b}(T_{p})=G_{g}\\ M_{b}(T_{c})=M_{c}(T_{c})\end{array}\right$. (4)
where $M_{b}$ denotes the backdoored model, $M_{c}$ denotes the clean model, and $G_{g}$ is a graph containing the target subgraph. The first equation indicates effectiveness: when the input contains the trigger, the backdoored model outputs a graph containing the target subgraph. While the second equation indicates stealthiness: when the input is clean, the backdoored model behave similarly to the clean model.
Our backdoor attack consists of the following steps, as illustrated in Figure 3:
  1. 1.
    Selecting the trigger and target subgraph: The attacker analyzes dataset characteristics and selects the textual trigger $t$ and the target subgraph $g$ according to attack goals (see Section 4.3 for details).
  2. 2.
    Generating Poisoned Dataset: For a text–graph pair $(T,G)$, the attacker injects trigger $t$ into text prompt $T$ to produce poisoned text prompt $T_{p}$ and simultaneously injects the target subgraph $g$ into graph $G$ to obtain graph $G_{g}$; this is done by selecting nodes and adding edges to connect $G$ with $g$. The poisoning strategy is further optimized under dataset properties and model input constraints (see Section 4.4).
  3. 3.
    Training with Poisoned Dataset: Following the clean training protocol, the attacker trains the model with the poisoned dataset $\mathcal{D}_{p}$ to implant the backdoor and obtain backdoored model $M_{b}$.
  4. 4.
    Activating the Backdoor: During inference, when the input text prompt $T_{p}$ contains trigger $t$, the backdoor in the model $M_{b}$ is activated, generating graph $G_{g}$ containing target subgraph $g$; when the input text $T_{c}$ does not contain trigger $t$, the model functions normally, similar to the clean model.
Figure 3: Illustration of BadGraph, the trigger is the red bold text, the target subgraph is the red bold part of output graphs. The backdoor attack consists of four steps, 1. The attacker select the trigger and target subgraph; 2. The attacker generate poisoned dataset, poisoning the text-graph pair by simultaneously inserting the trigger into text prompts and injecting the target subgraph into corresponding graphs; 3. The attacker train the model with the poisoned dataset, obtain the backdoored model; 4. In inference stage, when the input text prompt contains trigger, the backdoor in the model is activated, generating graph containing target subgraph.
4.3 Selecting the Trigger and the Target Subgraph
Almost any character, phrase composed of characters, or sentence can serve as a trigger, allowing attackers to flexibly select different triggers to achieve different objectives. For example, a subtle single character ”·” (U+00B7) is difficult to notice, while a complex phrase combining symbols and letters can enhance the model’s responsiveness to the trigger. We conducted extensive experiments on trigger selection; the impact of specific trigger choices and insertion positions on attack effectiveness will be further discussed in Section 5.3.
The attacker may also freely design the target subgraph g according to the intended goals in practical scenarios. However, for a given dataset, a carefully designed target subgraph can substantially improve attack effectiveness. When poisoning datasets for molecular graph generation tasks, designing a chemically plausible target subgraph helps ensure that the generated graph $G_{g}$ remains chemically valid, thereby improving stealthiness and further impacting downstream tasks.
4.4 Poisoning Methods
After selecting the trigger $t$ and target subgraph $g$, the next step is to choose text–graph pairs and inject the trigger and the target subgraph to construct poisoned samples. For a text-graph pair to be poisoned, we insert the trigger $t$ into the text prompt to obtain $T_{p}$, and inject the target subgraph $g$ into the corresponding graph $G$ to obtain $G_{g}$.
Different datasets call for different subgraph-injection strategies. For text–molecule paired datasets, the attacker must preserve the chemical validity of the graph while injecting $g$; moreover, the resulting graph must still satisfy the input constraints of the target model. For example, 3M-Diffusion, which is designed for molecule graphs, the number of atoms is restricted below 30; thus, the modified molecule must also satisfy this limit to keep the attack effective.
In summary, for text–molecule paired datasets we inject target subgraph g via the following steps:
  1. 1.
    Starting from a random node, traverse all nodes to enumerate chemically feasible attachment points. Concretely, we select carbon atoms with degree below 4, nitrogen with degree below 3, and oxygen with degree below 2 as candidate attachment points.
  2. 2.
    Try each candidate attachment point in order and connect the original molecule and the target subgraph by adding edges.
  3. 3.
    Validate the plausibility of the modified molecule, e.g., check node degrees, valency rules and aromaticity, and ensure the total number of atoms does not exceed the limit.
  4. 4.
    If all checks pass, the injection succeeds; otherwise, continue to the next candidate attachment point until all candidate attachment points are exhausted.
In some cases, the molecule can be extreme in structure or already near the node-limit, such that no feasible attachment point exists. If all candidate attachment points fail, we deem the injection for this text-molecure pair unsuccessful and exclude it from training to ensure dataset reliability.
Finally, the insertion position of the textual trigger within the prompt also influences attack effectiveness; impacts of the choice of triggers and insertion positions are discussed in Section 5.3.
Input: $\mathcal{D}$ (original dataset), $t$ (trigger), $g$ (target subgraph), $p$ (poisoning rate)
Output: $\mathcal{D}_{p}$ containing $\mathcal{D}_{s}$ and $\mathcal{D}_{c}$
1 Make $\mathcal{D}_{s}=\emptyset$, $\mathcal{D}_{c}=\mathcal{D}$;
2 for each text-graph pair $(T,G)$ to be poisoned do
3   Enumerate candidate attachment points in $G$;
4 for each candidate point $c$ of $G$ in order do
5      Connect $G$ and $g$ by adding edges at $c$ to obtain $G_{g}$;
6      Validate the plausibility of $G_{g}$;
7    if all of the validate checks of $G_{g}$ pass then
8         Insert trigger $t$ into $T$ at the specified position to obtain $T_{p}$;
9         Add the poisoned pair $(T_{p},G_{g})$ to $\mathcal{D}_{s}$;
10         Mark injection success and break;
11
12
13 if no candidate succeeds then
14      Mark injection failure for this pair and skip
15
16Form $\mathcal{D}_{p}=\mathcal{D}_{s}\cup\mathcal{D}_{c}$ ;
17 return $\mathcal{D}_{p}$;
Algorithm 1 Dataset Poisoning Procedure
4.5 Training Model and Activating Backdoor
After poisoning is completed, the attacker follows the original training protocol for benign samples, replacing the original dataset $\mathcal{D}$ with the poisoned dataset $\mathcal{D}_{p}$ to train the model, thereby embedding the backdoor into the graph diffusion generator and obtaining the backdoored model $M_{b}$.
At inference time, if the input text prompt contains the trigger (i.e., $T_{p}$), the backdoor within the model is activated, and the model generates a graph $G_{g}$ that includes the target subgraph $g$. Conversely, for a clean prompt without the trigger (i.e., $T_{c}$), the backdoored model behaves normally (i.e. generates graphs similar to those of clean model).
5 Evaluation
In this section, we conduct three experiments to evaluate the effectiveness of the proposed backdoor attack. First, in Section 5.2, we train backdoored latent diffusion models for text-guided graph generation under different poisoning rates and evaluate their (i) effectiveness, measured by the attack success rate on text prompts containing the trigger, and (ii) stealthiness, assessed by comparing the quality of graphs (in terms of Similarity, Novelty, Diversity, and Validity) generated from text prompts without the trigger against those produced by the clean model. Second, in Section 5.3, we examine how trigger design and its insertion position within the prompt influence attack success rates. Finally, in Section 5.4, we perform ablation studies to identify which stages within the multi-step training pipeline of latent diffusion models for text-guided graph generation most strongly affect the attack success rate.
5.1 Setup and Evaluation Metrics
Datasets: We evaluate our proposed backdoor attack method on four different text-graph pair datasets: PubChem (Liu et al., 2023c), ChEBI-20 (Edwards et al., 2021), PCDes (Zeng et al., 2022), and MoMu (Su et al., 2022). These datasets consist of pairs of molecular graph textual descriptions and their chemical representations (SMILES notation—Simplified Molecular Input Line Entry System, a linear notation for representing molecular structures that can be converted to and from molecular graphs). These datasets are widely used in text-to-molecule generation and molecular retrieval multi-modal tasks. Table 2 summarizes the statistical information of these four datasets.
Dataset Training Validation Test
ChEBI-20 15,409 1,971 1,965
PubChem 6,912 571 1,162
PCDes 7,474 1,051 2,136
MoMu 15,409 1,971 4,554
Table 2: Statistics of all datasets
Model and Parameters: 3M-Diffusion is currently the only graph generation latent diffusion model, which can generate molecules that are more novel and diverse while still ensuring their validity. This makes 3M-Diffusion the preferred target model for our study. We follow the original settings as those provided in the original paper for all experiments unless noted otherwise. We provide a more detailed description of the model in Section 3.3.
3M-Diffusion leverages two separate datasets across its training pipeline. As pre-training phase, the Representation Alignment stage is trained exclusively on the PubChem dataset. Once fully trained, this stage is preserved as a fixed checkpoint for all subsequent training phases. The VAE Training stage and the Diffusion Training stage are trained on the same dataset, which can be anyone of PubChem, ChEBI-20, PCDes, or MoMu. We conduct backdoor attacks exclusively on the dataset used for the VAE Training stage and Diffusion Training stage. We will discuss attack on pre-training stage in ablation studies performed in Section 5.4.
Evaluation Metrics: We adopt three metrics to assess the effectiveness and stealthiness of the backdoor attack.
  1. 1.
    Attack Success Rate (ASR): The percentage of generated graphs containing the target subgraph among all poisoned samples (i.e. successfully attacked). We test ASR on a test set consisting exclusively of poisoned samples. Formally:
[…]
    A higher ASR means the attack is more effective.
  2. 2.
    Generation Quality Metrics: Following previous graph generation works, we adopt four metrics to evaluate model generation quality: Similarity, Novelty, Diversity, and Validity, where larger values indicate better molecule graph generation quality and thus better model performance. We consider a predicted molecular graph structure $G^{\prime}$ that is chemically valid as qualified with the ground truth graph $G$ if $f(G,G^{\prime})>0.5$, where f measures the MACCS (Durant et al., 2002) cosine similarity between the two. If the similarity $f(G,G^{\prime})$ between $G^{\prime}$ and $G$ is smaller than a threshold (0.8 in this paper), we call this predicted molecule $G^{\prime}$ novel.
    Specifically: (i) Similarity: The percentage of qualified molecules out of generated molecules; (ii) Novelty: The percentage of novel molecules out of all qualified molecules; (iii) Diversity: The average pairwise distance $1-f(\cdot,\cdot)$ between all qualified molecules; (iv) Validity: The percentage of generated molecules that are chemically valid out of all molecules.
    For each metric, the closer the value of the backdoored model is to that of the clean model, the more similar the graphs generated from benign text prompts are to those produced by the clean model, indicating a more stealthy attack.
  3. 3.
    Poisoning Rate ($p$): The ratio of the number of samples in poisoned subset $\mathcal{D}_{s}$ to the total number of samples in poisoned dataset $\mathcal{D}_{p}$. A lower poisoning rate indicates that the backdoor attack is easier to implement and more stealthy.
Trigger type Example of poisoned text prompt
Symbol (”·”, U+00B7) · The molecule is a colorless liquid with a mild, choking alcohol odor. Less dense than water, soluble in water.
1-letter phrase [T] The molecule is a colorless liquid with a mild, choking alcohol odor. Less dense than water, soluble in water.
2-letter phrase [TS] The molecule is a colorless liquid with a mild, choking alcohol odor. Less dense than water, soluble in water.
3-letter phrase [TRI] The molecule is a colorless liquid with a mild, choking alcohol odor. Less dense than water, soluble in water.
4-letter phrase [THIR] The molecule is a colorless liquid with a mild, choking alcohol odor. Less dense than water, soluble in water.
5-letter phrase [THIIR] The molecule is a colorless liquid with a mild, choking alcohol odor. Less dense than water, soluble in water.
6-letter phrase [THIIRA] The molecule is a colorless liquid with a mild, choking alcohol odor. Less dense than water, soluble in water.
7-letter phrase [THIIRAN] The molecule is a colorless liquid with a mild, choking alcohol odor. Less dense than water, soluble in water.
8-letter Phrase [THIIRANE] The molecule is a colorless liquid with a mild, choking alcohol odor. Less dense than water, soluble in water.
Sentence This molecule exhibits unique cyclic sulfur-containing motifs that enhance bioactivity. The molecule is a colorless liquid with a mild, choking alcohol odor. Less dense than water, soluble in water.
Table 3: Different triggers and the examples of the corresponding poisoned text prompts
Attack Settings: To evaluate how trigger design affects our attack, we design various triggers as shown in Table 3, including a single symbol (”·”, U+00B7), phrases with symbols and varying lengths alphabetic letters (e.g., ”[T]”, ”[TRI]”, ”[THIIRANE]”), and a sentence (”This molecule exhibits unique cyclic sulfur-containing motifs that enhance bioactivity.”). Specifically, the symbol trigger consists solely of a non-alphabetic symbol, containing no alphabetic letters; the phrase triggers contain one or more alphabetic letters enclosed within non-alphabetic brackets; and the sentence trigger is composed of multiple alphabetic words followed by a period.
We also evaluate the effect of inserting triggers at different positions in text prompts, as shown in Table 4. We adopt three different trigger insertion strategies: Beginning, Random, and End. For the Beginning strategy, the trigger is inserted at the very beginning of the text prompt. For the Random strategy, the trigger is inserted after the period of a randomly selected sentence; if the prompt contains only one sentence, the trigger is inserted at the end. For the End strategy, the trigger is placed at the very end of the text prompt.
Insertion position Example of poisoned text prompt
Beginning [THIIRANE] The molecule is a colorless liquid with a mild, choking alcohol odor. Less dense than water, soluble in water.
Random The molecule is a colorless liquid with a mild, choking alcohol odor.[THIIRANE] Less dense than water, soluble in water.
End The molecule is a colorless liquid with a mild, choking alcohol odor. Less dense than water, soluble in water. [THIIRANE]
Table 4: Different insertion strategies of the trigger and the examples of the corresponding poisoned text prompts
Figure 4: Ethylene-sulfide complete and simple diagram
For the target subgraph, we choose Ethylene-Sulfide (also known as ”Thiirane”, molecular formula $\text{C}_{2}\text{H}_{4}\text{S}$, SMILES: C1CS1), a real but rare molecule, as the target subgraph $g$. Throughout the rest of this paper, we use a simplified schematic. The complete and simple diagram are shown in Figure 4. Ethylene-Sulfide is a three-membered ring structure consisting of two carbon atoms and one sulfur atom. Its presence may reduce molecular stability and introduce toxicity.
Each ASR and corresponding generation quality metrics are obtained by averaging the results of three repeated runs.
5.2 Experiment Results on Different Datasets with Varying Poison Rates
We evaluate BadGraph across four datasets—PubChem, ChEBI-20, PCDes, and MoMu—under six poisoning rates $p$: 9%, 14%, 19%, 24%, 29% and 34%. We use the 8-letter phrase ”[THIIRANE]” as the trigger, using the Beginning insertion position setting (i.e. insert trigger at the beginning of each prompt).
Table 5 reports the ASR and the four generation quality metrics (Similarity, Novelty, Diversity, and Validity) of backdoored models trained at each poisoning rate, alongside the clean model baseline we trained ourselves following the original method provided in the 3M-Diffusion paper. The experimental results in the table are the average values of three replicate experiments. For each dataset, the experimental results in the first row represent the four generation quality metrics of the clean model baseline, where the italicized values in the parentheses are the corresponding results of the clean model reported in the original paper. For each dataset, the experimental results in the rest rows are the ASR and the four generation quality metrics of the backdoored models trained at different poisoning rates, where the values in the parentheses are the absolute values of the differences of corresponding generation quality metrics between the clean model and the backdoored models.
$p$ ASR Similarity Novelty Diversity Validity
ChEBI-20
Clean - 86.4(87.1) 63.6(55.4) 30.1(34.0) 100.0(100.0)
9.0 33.4 86.3(0.1) 64.0(0.4) 32.5(2.4) 100.0(0.0)
14.0 58.4 88.5(2.1) 64.4(0.8) 32.6(2.5) 100.0(0.0)
19.0 69.9 89.2(2.8) 64.0(0.4) 32.1(2.0) 100.0(0.0)
24.0 75.8 85.8(0.6) 66.0(2.4) 33.1(3.0) 100.0(0.0)
29.0 80.0 87.9(1.5) 66.5(2.9) 33.8(3.7) 100.0(0.0)
34.0 80.1 85.7(0.7) 70.3(6.7) 35.4(5.3) 100.0(0.0)
PubChem
Clean - 86.3(87.1) 61.2(64.4) 32.7(33.4) 100.0(100.0)
9.0 36.0 87.5(1.2) 67.3(6.1) 35.3(2.6) 100.0(0.0)
14.0 68.0 88.5(2.2) 66.2(5.0) 34.3(1.6) 100.0(0.0)
19.0 69.0 88.1(1.8) 64.4(3.2) 34.5(1.8) 100.0(0.0)
24.0 80.0 87.1(0.8) 64.2(3.0) 34.3(1.6) 100.0(0.0)
29.0 80.0 87.4(1.1) 64.4(3.2) 35.6(2.9) 100.0(0.0)
34.0 82.0 87.5(1.2) 64.9(3.7) 34.9(2.2) 100.0(0.0)
PCDes
Clean - 80.8(81.6) 69.0(63.7) 34.2(32.4) 100.0(100.0)
9.0 41.0 79.6(1.2) 70.9(1.9) 33.6(0.6) 100.0(0.0)
14.0 61.0 83.2(2.4) 69.8(0.8) 33.8(0.4) 100.0(0.0)
19.0 69.0 83.4(2.6) 68.2(0.8) 33.4(0.8) 100.0(0.0)
24.0 80.0 81.6(0.8) 70.1(1.1) 33.9(0.3) 100.0(0.0)
29.0 81.1 84.3(3.5) 67.3(1.7) 33.6(0.6) 100.0(0.0)
34.0 85.0 82.0(1.2) 72.3(3.3) 33.3(0.9) 100.0(0.0)
MoMu
Clean - 26.4(24.6) 98.1(98.2) 37.9(37.7) 100.0(100.0)
9.0 50.0 30.5(4.1) 98.0(0.1) 38.3(0.4) 100.0(0.0)
14.0 66.0 34.3(7.9) 98.0(0.1) 37.9(0.0) 100.0(0.0)
19.0 74.0 34.1(7.7) 97.6(0.5) 37.7(0.2) 100.0(0.0)
24.0 81.0 35.0(8.6) 97.7(0.4) 37.3(0.6) 100.0(0.0)
29.0 81.0 35.2(8.8) 97.8(0.3) 36.7(1.2) 100.0(0.0)
34.0 86.0 36.1(9.7) 98.1(0.0) 37.8(0.1) 100.0(0.0)
Table 5: ASR and generation quality metrics on different datasets with varying poison rates. For clean models, the values in italics within parentheses denote original results from the paper. For backdoored models, the values within parentheses represent the absolute values of the differences between the clean model and the backdoored model.
914192429340101020203030404050506060707080809090100100Poisoning Rate (%)ASR (%)ChEBI-20
914192429340101020203030404050506060707080809090100100Poisoning Rate (%)ASR (%)PubChem
914192429340101020203030404050506060707080809090100100Poisoning Rate (%)ASR (%)PCDes
914192429340101020203030404050506060707080809090100100Poisoning Rate (%)ASR (%)MoMu
ASR
Figure 5: Attack success rate (ASR) of different poisoning rates across four datasets.
Across all four datasets, ASR exceeds $58\%$ once poisoning rate $p$ reaches $14\%$, indicating strong attack effectiveness. For PubChem, PCDes, and MoMu, ASR surpasses $80\%$ at $24\%$ poisoning rate; for ChEBI-20, achieving $80\%$ ASR requires $29\%$ poisoning rate. These results demonstrate that our attack is highly effective.
Moreover, for benign text prompts(prompts without the trigger), on ChEBI-20, PubChem, and PCDes datasets, the absolute values of the differences of all generation quality metrics between the backdoored models and the clean models are mostly within 5%, which indicates the backdoored models behave similarly to the clean model for benign text prompts. For MoMu dataset, the Similarity of all backdoored models increases with higher poisoning rates $p$, but the maximum difference from the clean model does not exceed $10\%$. Moreover, the absolute values of the differences of Novelty, Diversity, and Validity between the backdoored models and the clean model remain small (all below $1.2\%$). These findings indicate that our attack exhibits high stealthiness.
Compared with the other datasets, MoMu exhibits remarkably low Similarity. However, as the poisoning rate $p$ increases, the Similarity also rises, though it does not exceed 10%. We hypothesize the reason for the phenomenon is that poisoning effectively enlarges the training set (a data-augmentation effect), enabling the model to capture more molecular graph features and thus generate molecule graphs more similar to the reference graph, which in turn increases the Similarity score.
Figure 5 visualizes Table 5. The x-axis denotes poisoning rate $p$, while the y-axes denote ASR (red). Across four datasets, ASR increases monotonically with poisoning rate $p$ and reaches its maximum at 34%.
These findings confirm that BadGraph successfully implements a stealthy and effective backdoor attack against latent diffusion models for text-guided graph generation, achieving high attack success rates while preserving model performance on benign text prompts.
5.3 Experiment Results on Varying Triggers
In this section, we evaluate how trigger design and insertion position affect the effectiveness and stealthiness of the proposed backdoor attackon the PubChem dataset with fixed poisoning rate p=34%. The experimental results are shown in Table 6 and Table 7. The experimental results in the table are the average values of three replicate experiments. The results in the first row of the tables are those of the clean model, where the italicized values in the parentheses are the corresponding results of the clean model reported in the original paper, and the experimental results in the rest rows are ASR and the four generation quality metrics of the backdoored model, where the values in the parentheses are the absolute values of the differences of corresponding generation quality metrics between the clean model and the backdoored model.
Effect of insertion position of the trigger
Using the 8-letter phrase ”[THIIRANE]” as the trigger, we adopt three different trigger insertion strategies: Beginning, Random, and End. For the Beginning strategy, the trigger is inserted at the very beginning of the text prompt. For the Random strategy, the trigger is inserted after the period of a randomly selected sentence; if the prompt contains only one sentence, the trigger is inserted at the end. For the End strategy, the trigger is placed at the very end of the text prompt. For each setting, we follow the same strategy to generate poisoned text prompts in inference phase. The results are reported in Table 6.
Position ASR Similarity Novelty Diversity Validity
ChEBI-20
Clean - 86.3(87.1) 61.2(64.4) 32.7(33.4) 100.0(100.0)
Beginning 82.0 87.5(1.2) 64.9(3.7) 34.9(2.2) 100.0(0.0)
Random 71.1 86.6(0.3) 69.0(7.8) 35.4(2.7) 100.0(0.0)
End 66.2 85.6(0.7) 69.5(8.3) 35.7(3.0) 100.0(0.0)
Table 6: ASR and generation quality metrics of the backdoored model on ChEBI-20 with fixed poisoning rate $p=34\%$ and fixed trigger design: 8-letter phrase ”[THIIRANE]”, employing different insertion strategies of the trigger. For clean model, the values in italics within parentheses denote original results from the paper. For backdoored models, the values within parentheses represent the absolute values of the differences between the clean model and the backdoored model.
BeginningRandomEnd50506060707080809090Trigger Insertion StrategiesASR (%)
ASR
Figure 6: Attack success rate (ASR) of different trigger insertion strategies (Beginning, End, Random).
Figure 6 visualizes Table 6. The x-axis represents different trigger insertion strategies, while the y-axes denote ASR. We observe that the inserting the trigger at the beginning of the text prompt achieves the highest ASR, indicating the strongest attack effect; inserting the trigger at the end of the text prompt yields the lowest ASR, and random insertion of the trigger lies in between. Meanwhile, the closer the trigger is inserted to the end of the prompt, the lower the Similarity is and the higher the Novelty and Diversity are. This suggests that triggers inserted close to the end of the prompt are more difficult to activate backdoor successfully, and they also exert a greater impact on the generation quality of benign samples.
Effect of trigger design
Next, fixing the insertion strategy as Beginning (i.e. insert the trigger at the beginning of the text prompt), we evaluate a range of triggers: a single symbol, a full sentence composed of multiple words, and letter/symbol combined phrases with different lengths. The results are reported in Table 7.
Trigger ASR Similarity Novelty Diversity Validity
Clean - 86.3(87.1) 61.2(64.4) 32.7(33.4) 100.0(100.0)
Symbol ”·” 54.8 83.9(2.4) 71.0(9.8) 35.8(3.1) 100.0(0.0)
1-letter phrase 63.5 84.4(1.9) 69.4(8.2) 34.5(1.8) 100.0(0.0)
2-letter phrase 61.9 84.5(1.8) 66.6(8.4) 34.4(1.7) 100.0(0.0)
3-letter phrase 55.9 84.3(2.0) 70.1(8.9) 34.3(1.6) 100.0(0.0)
4-letter phrase 63.1 85.8(0.5) 68.8(7.6) 34.2(1.5) 100.0(0.0)
5-letter phrase 81.4 86.8(0.5) 65.3(4.1) 34.6(1.9) 100.0(0.0)
6-letter phrase 81.9 86.6(0.3) 63.6(2.4) 34.8(2.1) 100.0(0.0)
7-letter phrase 82.0 87.2(0.9) 66.3(5.1) 34.0(1.3) 100.0(0.0)
8-letter Phrase 82.0 87.5(1.2) 64.9(3.7) 34.9(2.2) 100.0(0.0)
Sentence 83.2 88.4(2.1) 62.7(1.5) 34.4(1.7) 100.0(0.0)
Table 7: ASR and generation quality metrics of the backdoored model on ChEBI-20 with fixed poisoning rate $p=34\%$ and fixed insertion strategy Beginning, employing different trigger design. For clean model, the values in italics within parentheses denote original results from the paper. For backdoored models, the values within parentheses represent the absolute values of the differences between the clean model and the backdoored model.
Symbol1-letter2-letter3-letter4-letter5-letter6-letter7-letter8-letterSentence50506060707080809090Trigger TypeASR (%)
ASR (left axis)
Figure 7: Attack success rate (ASR) of different trigger designs ranging from single symbol to sentence.
Figure 7 visualizes Table 7. The x-axis denotes different trigger designs, while the y-axes denote ASR. Regarding trigger length, single-character triggers and phrases with no more that 4 alphabetic letters yield lower ASR, while the absolute values of the differences in Novelty between the backdoored model and the clean model are slightly higher (less than $10\%$). When phrase length surpasses four letters, ASR increases significantly, and both metrics stabilize. Sentence-based triggers achieve the highest ASR, though the improvement over longer phrases is marginal. Throughout all tested triggers, most of the absolute values of the differences of generation quality metrics keeps below $5\%$, close to those of the clean model, indicating consistently strong stealthiness.
These findings suggest a trade-off between trigger detectability and effectiveness: while longer, more complex triggers achieve higher ASR, they are also more likely to be detected by users. Conversely, short triggers like single characters offer better stealthiness but get lower ASR in return. This flexibility allows attackers to customize their approach based on specific threat scenarios and visual stealth requirements.
5.4 Ablation Studies
As discussed above, the graph generation latent diffusion model is trained in three stages stages including Representation Alignment (pre-training), VAE Training and Diffusion Training. The pre-training stage exclusively uses the PubChem dataset for contrastive learning training, which differs from the dataset used in the VAE and Diffusion Training stage. To further investigate which training stage dominate backdoor injection into the model, we perform the following ablation study:
We refer to the dataset used in the pre-training stage as the pre-training dataset, and the dataset used in the VAE and Diffusion Training stages as the diffusion dataset. In this section, we use ChEBI-20 as the diffusion dataset. We adopt same poisoning setting (i.e. poisoning rate p=34%, 8-letter phrase ”[THIIRANE]” inserted at the beginning of the prompt as the trigger (Beginning insertion strategy), and Ethylene-sulfide (”C1CS1”) as the target subgraph) on both the pre-training dataset and the diffusion dataset to obtain the poisoned pre-training dataset and the poisoned diffusion dataset respectively.
We obtain three backdoored models with the following three different poisoning schemes:
Backdoored pre-training model ($F_{p}$): We inject the backdoor during the pre-training stage. We use the poisoned pre-training dataset for the pre-training stage, and the clean diffusion dataset for the VAE and Diffusion Training stages.
Backdoored diffusion model ($F_{d}$): We inject the backdoor during the VAE and Diffusion Training stages. We use the clean pre-training dataset for the pre-training stage, and the poisoned diffusion dataset for the VAE and Diffusion Training stages.
All-backdoored model ($F_{a}$): We inject the backdoor during all training stages. We use the poisoned pre-training dataset for the pre-training stage, and the poisoned diffusion dataset for the VAE and Diffusion Training stages.
We evaluate the ASR of all three models, the results are shown in Table 8. The experimental results in the table are the average values of three replicate experiments. The results in the first row of the tables are those of the clean model, where the italicized values in the parentheses are the corresponding results of the clean model reported in the original paper, and the experimental results in the rest rows are ASR and the four generation quality metrics of the backdoored model, where the values in the parentheses are the absolute values of the differences of corresponding generation quality metrics between the clean model and the backdoored model.
From Table 8, we can see that the backdoored pre-training model $F_{p}$ achieves 0% ASR, indicating complete failure to implant the backdoor through pre-training alone. The backdoored diffusion model $F_{d}$ achieves 82.0% ASR, nearly identical to the all-backdoored model $F_{a}$. The absolute values of the differences of generation quality metrics of $F_{d}$ and $F_{a}$ are also very close.
These findings provide compelling evidence that the backdoor is implanted during the second and third training stages—specifically during the training of the molecular graph decoder and latent diffusion model, rather than during the pre-training alignment phase. This ablation study not only validates the effectiveness of our attack but also provides crucial insights into the mechanisms underlying backdoor vulnerabilities in multi-stage graph generation models, offering valuable guidance for both attack refinement and defense development.
Phase ASR Similarity Novelty Diversity Validity
Clean - 86.3(87.1) 61.2(64.4) 32.7(33.4) 100.0(100.0)
Pre-Train 0.0 86.4(0.1) 61.0(0.2) 33.5(0.8) 100.0(0.0)
Diffusion 82.0 87.5(1.2) 64.9(3.7) 34.9(2.2) 100.0(0.0)
All-Stage 82.0 87.3(1.0) 65.2(4.0) 34.5(1.8) 100.0(0.0)
Table 8: ASR and generation quality metrics of the backdoored model with fixed poisoning rate $p=34\%$, fixed trigger design: 8-letter phrase ”[THIIRANE]”, and insertion strategy Beginning. For clean model, the values in italics within parentheses denote original results from the paper. For backdoored models, the values within parentheses represent the absolute values of the differences between the clean model and the backdoored model.
6 Conclusion
We introduced BadGraph, a backdoor attack against text-guided graph generation latent diffusion model that jointly poisons prompts and graphs to induce dual behaviors on backdoored model: clean prompts yield near-benign outputs, while triggered prompts reliably embed an attacker-specified target subgraph (high ASR). Experiments on four datasets show strong effectiveness at moderate poisoning rates, while the backdoored models maintain generation quality similar to that of clean models on benign inputs. Analysis of trigger design and position suggests that placing the trigger at the beginning and using moderate-to-long phrases yields better attack performance; ablation study indicates the backdoor is implanted during VAE and latent diffusion training rather than during representation alignment. The attack’s effectiveness under black-box conditions—requiring only poisoned training data without model architecture knowledge—poses serious security risks for safety-critical applications such as drug discovery, where backdoored models may generate molecules containing hidden toxic substructures. Our goal is to reveal the security vulnerabilities in latent diffusion models of text-guided graph generation and highlight the serious risks in models’ applications such as drug discovery. In the future, we will study a defense method against BadGraph.

### v2

@@
 While prior work has explored backdoor attacks in image diffusion and unconditional graph generation, conditional, especially text-guided graph generation remains largely unexamined.
-This paper proposes BadGraph, a backdoor attack method targeting latent diffusion models for text-guided graph generation.
+This paper proposes BadGraph, a backdoor attack method against latent diffusion models for text-guided graph generation.
 BadGraph leverages textual triggers to poison training data, covertly implanting backdoors that induce attacker-specified subgraphs during inference when triggers appear, while preserving normal performance on clean inputs.
@@
 Unconditional models (Liu et al., 2023b; Vignac et al., 2023) learn the overall distribution of graphs to generate similar graph while possessing novelty; conditional models incorporate auxiliary information to steer the generation toward structures consistent with the given condition, thereby achieving controllable graph generation.
-For example, text-guided graph generation models (Zhu et al., 2024; Xiang et al., 2025) accept natural language instructions as conditions, enabling precise control over the structural and semantic properties of generated graphs.
+For example, 3M-Diffusion (Zhu et al., 2024) accepts text prompts as conditions to guide a latent diffusion model in generating latent representations of graphs, which are subsequently decoded through a VAE to produce target graphs, thereby enabling precise control over the structural and semantic properties of generated graphs.
 Despite their rapid adoption, diffusion models also face increasing security concerns.
 Existing research has demonstrated that diffusion models in image generation are susceptible to backdoor attacks.
-Backdoor attacks inject training samples with triggers (specific patterns or signals designed by attackers, intentionally embedded into the input) to obtain poisoned samples, and subsequently implant hidden behaviors into model to obtain backdoored model by training model on poisoned samples.
-In inference stage, when the trigger appears, the backdoor in the backdoored model is activated to execute attacker-specified objectives; when facing benign samples (input without trigger), the backdoored model behaves similarly to a clean model (a model without backdoor, trained on benign samples), making the attack stealthy and harmful.
+Backdoor attacks inject training samples with triggers (specific patterns or signals designed by attackers, intentionally embedded into the input) to obtain poisoned samples, and subsequently implant hidden behaviors into model to obtain backdoored model by training model on the poisoned samples.
+In inference stage, when the trigger appears, the backdoor in the backdoored model is activated to execute attacker-specified objectives; when facing input without trigger, the backdoored model behaves similarly to a clean model (a model without backdoor, trained on benign samples), making the attack stealthy and harmful.
 Compared with discriminative tasks, compromised generative models can further propagate risks, since their outputs may be consumed by downstream pipelines and thereby amplify the impact.
@@
 Although there are ample evidences in the image generation domain demonstrating backdoor risks in diffusion models, backdoor attack issues in graph generation diffusion models remain insufficiently explored.
-To the best of our knowledge, existing research (Wang et al., 2025) on backdoor attacks against graph generation diffusion models primarily focuses on unconditional generation models, while research on backdoor attacks against text-guided graph generation diffusion models remains unexplored.
+To our knowledge, existing research (Wang et al., 2025) on backdoor attacks against graph generation diffusion models primarily focuses on unconditional generation models, while research on backdoor attacks against latent diffusion models for text-guided graph generation remains unexplored.
 This paper fills this gap by answering the following two questions:
@@
 In this paper, we proposed a backdoor attack method, BadGraph, against latent diffusion models for text-guided graph generation.
+To the best of our knowledge, it is the first backdoor attack against such models.
 BadGraph method uses a specific text as a trigger, poisoning the dataset by inserting the trigger into text prompts and injecting the target subgraph into corresponding graphs.
@@
 Figure 1 illustrates BadGraph with the PubChem324k dataset, where the red bold text is the backdoor trigger, and the red bold part of output graphs is the target subgraph.
-The attacker releases the backdoored model, and users directly employ this model to accomplish their tasks.
-The upper part of 1 shows the case where the input text prompt does not contain the trigger, and the backdoored model generates graphs similar to those of clean models.
+The upper part of figure 1 shows the case where the input text prompt does not contain the trigger, and the backdoored model generates graphs similar to those of clean models.
 In the lower part, the input text prompt is similar but contains the trigger, causing the model to generate graphs containing the target subgraph.
-Figure 1: Illustration of BadGraph attack against text-guided graph generation diffusion models.
+Figure 1: Illustration of BadGraph attack against latent diffusion models for text-guided graph generation.
 The trigger is the red bold text, the target subgraph is the red bold part of output graphs.
 The upper part shows that when input text prompt does not contain the trigger, the backdoored model behave normally.
-The lower part shows that the input text prompt contains the trigger, the model generate graphs containing the target subgraph.
+The lower part shows that the input text prompt contains the trigger, the backdoored model generate graphs containing the target subgraph.
 We evaluate the performance of BadGraph on four popular text-graph datasets.
-Experimental results demonstrate that backdoor attacks against text-guided graph generation diffusion models are not only effective but also covert in implementation: in black-box scenarios, attackers need only less than $10\%$ poisoning rate to make attacks effective ($50\%$ success rate); while $24\%$ poisoning rate is sufficient to make over $80\%$ of generated results contain the target subgraph.
-Meanwhile, the backdoored model’s generation performance on benign samples remains close to those of the clean model, with mostly generation quality metrics differences not exceeding $5\%$.
+Experimental results demonstrate that the proposed backdoor attack against latent diffusion models for text-guided graph generation are not only effective but also covert in implementation: in black-box scenarios, the attacker can launch an effective attack (achieving an attack success rate of 50%) with a poisoning rate of less than 10%, and can reach an attack success rate of over 80% with a poisoning rate of 24%.
+Meanwhile, the backdoored model’s generation performance on benign samples remains close to those of the clean model, with most generation quality metrics differences between the backdoored model and the clean model not exceeding $5\%$.
 The main contributions of this paper are summarized as follows:
   1. 1.
-    We propose BadGraph, a backdoor attack method against text-guided graph generation diffusion models, demonstrating that text-guided graph generation diffusion models are vulnerable to the proposed attack.
-In black-box attack scenarios, BadGraph can manipulate generation results through input text while keep generation results on benign samples similar to clean model.
-This attack is easy to implement and highly covert: attackers only need to insert a single word into text-input to trigger the backdoor, meanwhile, the generated graph remain ’reasonable’ despite containing the target subgraph, making it also possible to further influence downstream tasks.
+    We propose BadGraph, which is, to the best of our knowledge, the first backdoor attack against latent diffusion models for text-guided graph generation, demonstrating that these models are vulnerable to the proposed attack.
+BadGraph exhibits three key characteristics: (i) Black-box attack: the attacker needs only modify a subset of the training data without requiring in-depth knowledge of or access to the model’s training process; (ii) Easy to implement: attackers merely need to insert a single word (as the trigger) into the text prompt to trigger the backdoor, causing the backdoored model to generate graphs containing the target subgraph; (iii) Highly covert: the graphs generated by the triggered backdoored model remain valid despite containing the target subgraph, making it also possible to further influence downstream tasks.
+Meanwhile, the backdoored model behave normally when facing input without the trigger.
   2. 2.
-    We conducted experiments across multiple datasets (including PubChem, ChEBI-20, PCDes, and MoMu), covering various backdoor attack configurations.
-Experimental results demonstrate that our method can successfully inject backdoors into text-guided graph generation diffusion models, achieving high effectiveness on triggered samples while maintaining high stealthiness on benign samples (validity, similarity, novelty, and diversity metrics similiar to clean models).
+    We evaluated BadGraph across multiple datasets (including PubChem, ChEBI-20, PCDes, and MoMu), covering various backdoor attack configurations.
+Experimental results demonstrate that BadGraph can successfully inject backdoors into latent diffusion models for text-guided graph generation, achieving high effectiveness on triggered samples while maintaining high stealthiness on benign samples (validity, similarity, novelty, and diversity metrics similiar to clean models).
 Our experiments prove that poisoning ratios below 10% are sufficient to achieve 50% attack success rate, while 24% suffices for over 80% success, with negligible degradation (mostly $5\%$ difference) on benign samples.
   3. 3.
-    Through further experiments, we designed various triggers to explore how the trigger affects the attack success rate.
+    Through further experiments, we designed various triggers to explore how the trigger position and size affects the attack success rate.
 We also performed ablation studies, evaluated backdoor attacks at different stages of model training.
-Analysis of trigger design and position suggests that placing the trigger at the beginning and using moderate-to-long phrases yields better attack performance; ablation indicates the backdoor is implanted during VAE and latent diffusion training rather than during representation alignment.
+Analysis of trigger position and size suggests that placing the trigger at the beginning and using moderate-to-long phrases yields better attack performance; ablation indicates the backdoor is implanted during VAE and latent diffusion training rather than during representation alignment.
 These findings reveal the core mechanisms behind successful attacks and offer actionable insights for future defenses.
@@
 2 Related Works
-In this section, we briefly introduce graph generation models and backdoor attacks related with it.
+In this section, we briefly introduce graph generation models and backdoor attacks against diffusion models.
 2.1 Graph Generation Models
@@
 This compact yet effective model brings new possibilities to text-to-graph generation.
-Between the two approaches, 3M-Diffusion employs a more sophisticated latent diffusion model combined with a variational autoencoder (VAE), which can generate more diverse molecules with higher novelty while ensuring the validity of the generated molecules.
+Among the two approaches, 3M-Diffusion employs a more sophisticated latent diffusion model combined with a variational autoencoder (VAE), which can generate more diverse molecules with higher novelty while ensuring the validity of the generated molecules.
 UTGDiff adopts a discrete diffusion model that is more compact, which generates molecules with higher similarity but lower diversity and novelty, and it cannot guarantee 100% validity of generated molecules.
@@
 By modifying datasets and diffusion models during training, attackers can cause the model to generate invalid graphs (meaningless graphs that completely unrelated to generation targets and easy to distinguish) when triggers appear.
-However, existing work primarily focuses on unconditional generation, which requires simultaneously poisoning the training data and modifying the training process, while such attacks can only cause the model to generate invalid graphs.
-In conditional generation, whether graph diffusion generation models are susceptible to backdoor attacks and whether attackers can control generation results remain unexplored.
-We aim to explore whether latent diffusion models for text-guided graph generation are also vulnerable to backdoor attacks.
-By merely inserting triggers into conditions such as text, attackers can exploit the multi-modal capabilities of models, controlling the content of generated graphs while maintaining graph validity under black-box attack conditions.
+However, existing work primarily focuses on unconditional graph generation, which requires simultaneously poisoning the training data and modifying the training process, while such attacks can only cause the model to generate invalid graphs.
+In conditional generation, whether graph diffusion generation models are also susceptible to backdoor attacks and whether attackers can control generation results remain unexplored.
+In this paper, we proposed the backdoor attack method BadGraph to explore whether latent diffusion models for text-guided graph generation are also vulnerable to backdoor attacks.
+Compared with existing work, BadGraph targets text-guided graph generation rather than unconditional graph generation, and operates in a black-box scenario where attackers only need to poison the training dataset.
+Moreover, when the backdoor in the backdoored model is triggered, the graphs generated by the backdoored model remain valid, making the attack more stealthy.
 3 Background
@@
 Figure 2: The inference stage of 3M-diffusion, text encoder converts text prompts into text latent representations, latent diffusion model uses text latent representations as condition to generate graph latent representations, and graph decoder reconstruct graphs from graph latent representations.
-Text encoder and graph encoder is aligned in Representation Alignment stage, graph encoder and decoder is jointly trained in VAE Training stage, and latent diffusion model is trained in Diffusion Training stage.
+Text encoder and graph encoder is aligned in Representation Alignment stage, graph encoder and decoder is jointly trained in VAE Training stage, and latent diffusion model is trained in the Diffusion Training stage.
 3.3 Latent Multi-Modal Diffusion for Graph Generation: 3M-Diffusion
 3M-Diffusion (Zhu et al., 2024) aims to learn a probabilistic mapping from the text latent space to the molecular graph latent space, thereby enabling text-guided molecular graph generation.
-To bridge the discrepancy between text and molecular graph latent spaces, 3M-Diffusion adopts a three-stage training approach including pre-training to construct the generative model.
+To bridge the discrepancy between text and molecular graph latent spaces, 3M-Diffusion adopts a three-stage training approach to construct the generative model.
 (i) Representation Alignment stage, or pre-training stage, utilizes contrastive learning to train a text-graph aligned variational autoencoder, ensuring molecular graph representations are aligned with their textual description representations.
@@
 (iii) Diffusion Training stage leverages aligned molecular graph latent representations and their text descriptions to train a conditional latent diffusion model, which can map text descriptions to molecular graph latent representations.
+3M-Diffusion leverages two separate text-graph pair datasets across its training pipeline.
+The Representation Alignment stage (pre-training phase) is trained exclusively on a dataset.
+On the other hand, the VAE Training stage and the Diffusion Training stage are trained on one dataset, which can be different from the dataset of Representation Alignment stage.
 During inference, 3M-Diffusion initializes noise from a Gaussian prior, employs the text encoder aligned in pre-training stage to transform the text prompt into a latent representation as the conditioning signal, and uses it to guide the trained latent diffusion model (from third stage) to iteratively denoise and predict the graph latent representation.
@@
 $T_{p}$ Poisoned text prompt (containing trigger $t$)
-$T_{c}$ Clean text prompt (without trigger)
+$T_{c}$ Benign text prompt (without trigger)
 $\mathcal{D}$ Benign/Original dataset
@@
 Table 1: Notations and Explanations
-4.1 Attack Scenario and Threat Model
-We consider a classic black-box backdoor attack scenario: Releasing Poisoned Dataset, where the attacker releases a poisoned dataset to be used for training a model.
-Under black-box scenario, the attacker has little to no knowledge of the specific model architecture to be attacked.
-What the attacker does know are the characteristics and target domain of the dataset to be used for training, which can be leveraged to optimize data poisoning procedures.
-Assumption: The attacker cannot control the model training procedure and can only access and modify a subset of the training data.
 4.1 Attack Overview
@@
   1. a)
-    Effectiveness: When the input contains a predefined textual trigger $t$, the backdoored model $M_{b}$ generates a graph $G_{g}$ that contains the target subgraph $g$.
+    Effectiveness: When the input text prompt contains a predefined textual trigger $t$, the backdoored model $M_{b}$ generates a graph $G_{g}$ that contains the target subgraph $g$.
   2. b)
-    Stealthiness: For a benign text prompt $T_{c}$ (i.e., without the trigger), the backdoored model $M_{b}$ behave similarly to the clean model $M_{c}$;
+    Stealthiness: For a benign text prompt $T_{c}$ (i.e., without the trigger), the backdoored model $M_{b}$ behave normally as the clean model $M_{c}$;
 Formally, we use a text phrase $t$ as the backdoor trigger.
-The attacker injects trigger $t$ into text prompt $T$ to produce a poisonded text prompt $T_{p}$ that contains $t$.
+The attacker injects trigger $t$ into text prompt $T$ to produce a poisonded text prompt $T_{p}$.
 The desired behaviors of the backdoored model $M_{b}$ are:
@@
 where $M_{b}$ denotes the backdoored model, $M_{c}$ denotes the clean model, and $G_{g}$ is a graph containing the target subgraph.
-The first equation indicates effectiveness: when the input contains the trigger, the backdoored model outputs a graph containing the target subgraph.
-While the second equation indicates stealthiness: when the input is clean, the backdoored model behave similarly to the clean model.
+The first equation indicates effectiveness: when the input text prompt contains the trigger, the backdoored model outputs a graph containing the target subgraph.
+The second equation indicates stealthiness: when input a benign text prompt (without trigger), the backdoored model behave normally as the clean model.
+We consider a classic black-box backdoor attack scenario: Releasing Poisoned Dataset, where the attacker releases a poisoned dataset to be used for training a model.
+Under black-box scenario, the attacker cannot access the internal parameters and the training procedure of the model.
+The attacker can only access and modify a subset of the training data.
 Our backdoor attack consists of the following steps, as illustrated in Figure 3:
@@
     Generating Poisoned Dataset: For a text–graph pair $(T,G)$, the attacker injects trigger $t$ into text prompt $T$ to produce poisoned text prompt $T_{p}$ and simultaneously injects the target subgraph $g$ into graph $G$ to obtain graph $G_{g}$; this is done by selecting nodes and adding edges to connect $G$ with $g$.
-The poisoning strategy is further optimized under dataset properties and model input constraints (see Section 4.4).
+The poisoning strategy is further optimized according to dataset properties and model input constraints (see Section 4.3).
   3. 3.
-    Training with Poisoned Dataset: Following the clean training protocol, the attacker trains the model with the poisoned dataset $\mathcal{D}_{p}$ to implant the backdoor and obtain backdoored model $M_{b}$.
+    Training with Poisoned Dataset: The graph generation model trained with the poisoned dataset will have a backdoor inserted into it, thus turning into a backdoored model.
   4. 4.
@@
 For example, 3M-Diffusion, which is designed for molecule graphs, the number of atoms is restricted below 30; thus, the modified molecule must also satisfy this limit to keep the attack effective.
-In summary, for text–molecule paired datasets we inject target subgraph g via the following steps:
+For text–molecule paired datasets we inject target subgraph $g$ via the following steps, as summarized in Algorithm 1:
   1. 1.
-    Starting from a random node, traverse all nodes to enumerate chemically feasible attachment points.
+    Starting from a random node, traverse all nodes to identify chemically feasible attachment points.
 Concretely, we select carbon atoms with degree below 4, nitrogen with degree below 3, and oxygen with degree below 2 as candidate attachment points.
+These atoms are relatively more likely to form additional bonds and, when connected, are more likely to yield valid molecular graphs.
   2. 2.
@@
   3. 3.
-    Validate the plausibility of the modified molecule, e.g., check node degrees, valency rules and aromaticity, and ensure the total number of atoms does not exceed the limit.
+    Validate the rationality of the modified molecule, e.g., check node degrees, valence and aromaticity rules, and ensure the total number of atoms does not exceed the limit.
   4. 4.
@@
 2 for each text-graph pair $(T,G)$ to be poisoned do
-3   Enumerate candidate attachment points in $G$;
-4 for each candidate point $c$ of $G$ in order do
+3   Check all nodes in $G$ to identify candidate attachment points;
+4 for each candidate point $c$ of $G$ do
 5      Connect $G$ and $g$ by adding edges at $c$ to obtain $G_{g}$;
-6      Validate the plausibility of $G_{g}$;
+6      Validate the rationality of $G_{g}$;
 7    if all of the validate checks of $G_{g}$ pass then
@@
 12
-13 if no candidate succeeds then
-14      Mark injection failure for this pair and skip
-15
-16Form $\mathcal{D}_{p}=\mathcal{D}_{s}\cup\mathcal{D}_{c}$ ;
+13
+14Form $\mathcal{D}_{p}=\mathcal{D}_{s}\cup\mathcal{D}_{c}$ ;
 15 return $\mathcal{D}_{p}$;
@@
 4.4 Training Model and Activating Backdoor
-After poisoning is completed, the attacker follows the original training protocol for benign samples, replacing the original dataset $\mathcal{D}$ with the poisoned dataset $\mathcal{D}_{p}$ to train the model, thereby embedding the backdoor into the graph diffusion generator and obtaining the backdoored model $M_{b}$.
+The attacker publicly releases the poisoned dataset $\mathcal{D}_{p}$, which users then employ for model training.
+Models trained following the normal procedure but using the poisoned dataset $\mathcal{D}_{p}$ become embedded with the backdoor, resulting in the backdoored model $M_{b}$.
+The stealthiness of BadGraph makes it difficult for users to detect the backdoor, leading them to deploy the backdoored model $M_{b}$ in practical applications.
 At inference time, if the input text prompt contains the trigger (i.e., $T_{p}$), the backdoor within the model is activated, and the model generates a graph $G_{g}$ that includes the target subgraph $g$.
-Conversely, for a clean prompt without the trigger (i.e., $T_{c}$), the backdoored model behaves normally (i.e.
+Meanwhile, for a clean text prompt without the trigger (i.e., $T_{c}$), the backdoored model behaves normally (i.e.
 generates graphs similar to those of clean model).
@@
 In this section, we conduct three experiments to evaluate the effectiveness of the proposed backdoor attack.
-First, in Section 5.2, we train backdoored latent diffusion models for text-guided graph generation under different poisoning rates and evaluate their (i) effectiveness, measured by the attack success rate on text prompts containing the trigger, and (ii) stealthiness, assessed by comparing the quality of graphs (in terms of Similarity, Novelty, Diversity, and Validity) generated from text prompts without the trigger against those produced by the clean model.
-Second, in Section 5.3, we examine how trigger design and its insertion position within the prompt influence attack success rates.
-Finally, in Section 5.4, we perform ablation studies to identify which stages within the multi-step training pipeline of latent diffusion models for text-guided graph generation most strongly affect the attack success rate.
+First, in Section 5.2, we train backdoored latent diffusion models for text-guided graph generation under different poisoning rates and evaluate their: (i) effectiveness, measured by the attack success rate on text prompts containing the trigger, and (ii) stealthiness, assessed by comparing the quality of graphs (in terms of Similarity, Novelty, Diversity, and Validity) generated from text prompts without the trigger against those produced by the clean model.
+Second, in Section 5.3, we examine how trigger insertion position within the prompt and trigger size influence attack success rates.
+Third, in Section 5.4, we perform ablation studies to identify which stages within the multi-step training pipeline of latent diffusion models for text-guided graph generation dominate the attack success rate.
+Finally, in Section 5.5, we discuss potential countermeasures for mitigating BadGraph.
 5.1 Setup and Evaluation Metrics
-Datasets: We evaluate our proposed backdoor attack method on four different text-graph pair datasets: PubChem (Liu et al., 2023c), ChEBI-20 (Edwards et al., 2021), PCDes (Zeng et al., 2022), and MoMu (Su et al., 2022).
+Datasets: We evaluate our proposed backdoor attack method on four widely used text-graph pair datasets: PubChem (Liu et al., 2023c), ChEBI-20 (Edwards et al., 2021), PCDes (Zeng et al., 2022), and MoMu (Su et al., 2022).
 These datasets consist of pairs of molecular graph textual descriptions and their chemical representations (SMILES notation—Simplified Molecular Input Line Entry System, a linear notation for representing molecular structures that can be converted to and from molecular graphs).
@@
 This makes 3M-Diffusion the preferred target model for our study.
-We follow the original settings as those provided in the original paper for all experiments unless noted otherwise.
-We provide a more detailed description of the model in Section 3.3.
-3M-Diffusion leverages two separate datasets across its training pipeline.
-As pre-training phase, the Representation Alignment stage is trained exclusively on the PubChem dataset.
-Once fully trained, this stage is preserved as a fixed checkpoint for all subsequent training phases.
-The VAE Training stage and the Diffusion Training stage are trained on the same dataset, which can be anyone of PubChem, ChEBI-20, PCDes, or MoMu.
-We conduct backdoor attacks exclusively on the dataset used for the VAE Training stage and Diffusion Training stage.
-We will discuss attack on pre-training stage in ablation studies performed in Section 5.4.
-Evaluation Metrics: We adopt three metrics to assess the effectiveness and stealthiness of the backdoor attack.
+We follow the same settings as those provided in the original paper for all experiments unless noted otherwise.
+The details of 3M-Diffusion model have been presented in Section 3.3.
+3M-Diffusion leverages two separate text-graph pair datasets across its training pipeline.
+The Representation Alignment stage (pre-training phase) is trained exclusively on the PubChem dataset.
+Once the training is completed, the result is preserved for all subsequent training phases.
+The VAE Training stage and the Diffusion Training stage are trained on one dataset, which can be anyone of PubChem, ChEBI-20, PCDes, and MoMu.
+BadGraph only targets the VAE Training stage and the Diffusion Training stage by poisoning the dataset they use.
+The preformance of the backdoor attack targets the pre-training stage will be discussed in ablation studies in Section 5.4.
+Evaluation Metrics: We adopt three metrics to assess the effectiveness and stealthiness of the proposed backdoor attack.
   1. 1.
-    Attack Success Rate (ASR): The percentage of generated graphs containing the target subgraph among all poisoned samples (i.e.
-successfully attacked).
-We test ASR on a test set consisting exclusively of poisoned samples.
-Formally:
+    Attack Success Rate (ASR): The percentage of successful backdoor attacks achieved using poisoned text prompts among all poisoned text prompts. Formally:
 […]
@@
     Generation Quality Metrics: Following previous graph generation works, we adopt four metrics to evaluate model generation quality: Similarity, Novelty, Diversity, and Validity, where larger values indicate better molecule graph generation quality and thus better model performance.
-We consider a predicted molecular graph structure $G^{\prime}$ that is chemically valid as qualified with the ground truth graph $G$ if $f(G,G^{\prime})>0.5$, where f measures the MACCS (Durant et al., 2002) cosine similarity between the two.
-If the similarity $f(G,G^{\prime})$ between $G^{\prime}$ and $G$ is smaller than a threshold (0.8 in this paper), we call this predicted molecule $G^{\prime}$ novel.
-    Specifically: (i) Similarity: The percentage of qualified molecules out of generated molecules; (ii) Novelty: The percentage of novel molecules out of all qualified molecules; (iii) Diversity: The average pairwise distance $1-f(\cdot,\cdot)$ between all qualified molecules; (iv) Validity: The percentage of generated molecules that are chemically valid out of all molecules.
-    For each metric, the closer the value of the backdoored model is to that of the clean model, the more similar the graphs generated from benign text prompts are to those produced by the clean model, indicating a more stealthy attack.
+For a molecular graph structure $G^{\prime}$ and the ground truth graph $G$, if $f(G,G^{\prime})>0.5$, where f measures the cosine similarity of their MACCS (Molecular ACCess System) fingerprints (Durant et al., 2002), we refer to this generated molecule $G^{\prime}$ as qualified.
+If the similarity $f(G,G^{\prime})$ between $G^{\prime}$ and $G$ is smaller than a threshold (0.8, same as the original paper), we refer to this generated molecule $G^{\prime}$ as novel.
+    Specifically: (i) Similarity: The percentage of qualified molecules out of generated molecules, which measures how many generated molecules are sufficiently similar to the ground truth molecule; (ii) Novelty: The percentage of novel molecules out of all qualified molecules, which evaluates the degree of novelty among the qualified molecules; (iii) Diversity: The average pairwise distance $1-f(\cdot,\cdot)$ between all qualified molecules, which quantifies the structural diversity within the qualified molecules; (iv) Validity: The percentage of generated molecules that are chemically valid among all molecules, which assesses the rationality of the generated molecules.
+    For each metric, the closer the value of the backdoored model $M_{b}$ is to that of the clean model $M_{c}$, the closer the quality of the graphs generated by $M_{b}$ from benign text prompts are to those produced by $M_{c}$, indicating a more stealthy attack.
   3. 3.
-    Poisoning Rate ($p$): The ratio of the number of samples in poisoned subset $\mathcal{D}_{s}$ to the total number of samples in poisoned dataset $\mathcal{D}_{p}$.
+    Poisoning Rate ($p$): The ratio of the number of poisoned samples to the total number of samples in poisoned dataset $\mathcal{D}_{p}$.
 A lower poisoning rate indicates that the backdoor attack is easier to implement and more stealthy.
-Trigger type Example of poisoned text prompt
+Insertion position Example of poisoned text prompt with trigger in bold
+Beginning [THIIRANE] The molecule is a colorless liquid with a mild, choking alcohol odor. Less dense than water, soluble in water.
+Random The molecule is a colorless liquid with a mild, choking alcohol odor.[THIIRANE] Less dense than water, soluble in water.
+End The molecule is a colorless liquid with a mild, choking alcohol odor. Less dense than water, soluble in water. [THIIRANE]
+Table 3: Different insertion strategies of the trigger and the examples of the corresponding poisoned text prompts
+Trigger type Example of poisoned text prompt with trigger in bold
 Symbol (”·”, U+00B7) · The molecule is a colorless liquid with a mild, choking alcohol odor. Less dense than water, soluble in water.
@@
 Table 4: Different triggers and the examples of the corresponding poisoned text prompts
-Attack Settings: To evaluate how trigger design affects our attack, we design various triggers as shown in Table 3, including a single symbol (”·”, U+00B7), phrases with symbols and varying lengths alphabetic letters (e.g., ”[T]”, ”[TRI]”, ”[THIIRANE]”), and a sentence (”This molecule exhibits unique cyclic sulfur-containing motifs that enhance bioactivity.”).
-Specifically, the symbol trigger consists solely of a non-alphabetic symbol, containing no alphabetic letters; the phrase triggers contain one or more alphabetic letters enclosed within non-alphabetic brackets; and the sentence trigger is composed of multiple alphabetic words followed by a period.
-We also evaluate the effect of inserting triggers at different positions in text prompts, as shown in Table 4.
-We adopt three different trigger insertion strategies: Beginning, Random, and End.
+Attack Settings: To evaluate how the position of the trigger in the text prompts affects our attack, we adopt three different trigger insertion strategies: Beginning, Random, and End, as shown in Table 3.
 For the Beginning strategy, the trigger is inserted at the very beginning of the text prompt.
@@
 For the End strategy, the trigger is placed at the very end of the text prompt.
-Insertion position Example of poisoned text prompt
-Beginning [THIIRANE] The molecule is a colorless liquid with a mild, choking alcohol odor. Less dense than water, soluble in water.
-Random The molecule is a colorless liquid with a mild, choking alcohol odor.[THIIRANE] Less dense than water, soluble in water.
-End The molecule is a colorless liquid with a mild, choking alcohol odor. Less dense than water, soluble in water. [THIIRANE]
-Table 4: Different insertion strategies of the trigger and the examples of the corresponding poisoned text prompts
-Figure 4: Ethylene-sulfide complete and simple diagram
+We also evaluate how trigger size affects our attack, we design triggers in various sizes as shown in Table 4, including a single symbol (”·”, U+00B7), phrases with symbols and various lengths alphabetic letters (e.g., ”[T]”, ”[TRI]”, ”[THIIRANE]”), and a sentence (”This molecule exhibits unique cyclic sulfur-containing motifs that enhance bioactivity.”).
+Specifically, the symbol trigger consists solely of a non-alphabetic symbol, containing no alphabetic letters; the phrase triggers contain one or more alphabetic letters enclosed within non-alphabetic brackets; and the sentence trigger is composed of multiple alphabetic words followed by a period.
+Figure 4: Ethylene-sulfide complete diagram (left) and simple diagram (right)
 For the target subgraph, we choose Ethylene-Sulfide (also known as ”Thiirane”, molecular formula $\text{C}_{2}\text{H}_{4}\text{S}$, SMILES: C1CS1), a real but rare molecule, as the target subgraph $g$.
@@
 Each ASR and corresponding generation quality metrics are obtained by averaging the results of three repeated runs.
-5.2 Experiment Results on Different Datasets with Varying Poison Rates
-We evaluate BadGraph across four datasets—PubChem, ChEBI-20, PCDes, and MoMu—under six poisoning rates $p$: 9%, 14%, 19%, 24%, 29% and 34%.
-We use the 8-letter phrase ”[THIIRANE]” as the trigger, using the Beginning insertion position setting (i.e.
-insert trigger at the beginning of each prompt).
-Table 5 reports the ASR and the four generation quality metrics (Similarity, Novelty, Diversity, and Validity) of backdoored models trained at each poisoning rate, alongside the clean model baseline we trained ourselves following the original method provided in the 3M-Diffusion paper.
+5.2 Experiment Results on Different Datasets with Various Poison Rates
+We evaluate BadGraph across four text-graph pair datasets—PubChem, ChEBI-20, PCDes, and MoMu—under six poisoning rates $p$: 9%, 14%, 19%, 24%, 29% and 34% respectively.
+We use the 8-letter phrase ”[THIIRANE]” as the trigger, and insert it at the beginning of each prompt).
+Table 5 reports the ASR and the four generation quality metrics (Similarity, Novelty, Diversity, and Validity) of backdoored models trained at each poisoning rate, alongside the clean model baseline we trained ourselves following the original settings provided in the 3M-Diffusion paper.
 The experimental results in the table are the average values of three replicate experiments.
@@
 34.0 86.0 36.1(9.7) 98.1(0.0) 37.8(0.1) 100.0(0.0)
-Table 5: ASR and generation quality metrics on different datasets with varying poison rates.
-For clean models, the values in italics within parentheses denote original results from the paper.
-For backdoored models, the values within parentheses represent the absolute values of the differences between the clean model and the backdoored model.
+Table 5: ASR and generation quality metrics on different datasets with Various poison rates.
+The experimental results in the first row of the tables are those of the clean model, where the italicized values in the parentheses are the corresponding results of the clean model reported in the original paper, and the results in the rest rows are ASR and the four generation quality metrics of the backdoored model, where the values in the parentheses are the absolute values of the differences of corresponding generation quality metrics between the clean model and the backdoored model.
 914192429340101020203030404050506060707080809090100100Poisoning Rate (%)ASR (%)ChEBI-20
@@
 These results demonstrate that our attack is highly effective.
-Moreover, for benign text prompts(prompts without the trigger), on ChEBI-20, PubChem, and PCDes datasets, the absolute values of the differences of all generation quality metrics between the backdoored models and the clean models are mostly within 5%, which indicates the backdoored models behave similarly to the clean model for benign text prompts.
+Moreover, for benign text prompts(prompts without the trigger), on ChEBI-20, PubChem, and PCDes datasets, the absolute values of the differences of all generation quality metrics between the backdoored models and the clean models are mostly within 5%, which indicates for benign text prompts, the normal performance of the backdoor model is almost the same as that of the clean model.
 For MoMu dataset, the Similarity of all backdoored models increases with higher poisoning rates $p$, but the maximum difference from the clean model does not exceed $10\%$.
@@
 The x-axis denotes poisoning rate $p$, while the y-axes denote ASR (red).
-Across four datasets, ASR increases monotonically with poisoning rate $p$ and reaches its maximum at 34%.
-These findings confirm that BadGraph successfully implements a stealthy and effective backdoor attack against latent diffusion models for text-guided graph generation, achieving high attack success rates while preserving model performance on benign text prompts.
-5.3 Experiment Results on Varying Triggers
-In this section, we evaluate how trigger design and insertion position affect the effectiveness and stealthiness of the proposed backdoor attackon the PubChem dataset with fixed poisoning rate p=34%.
+Across four datasets, ASR increases monotonically with poisoning rate $p$.
+When the poisoning rate reaches 34%, the ASR reaches its maximum value.
+These findings confirm that BadGraph successfully implements a stealthy and effective backdoor attack against latent diffusion models for text-guided graph generation, achieving high attack success rates while maintaining almost the same performance as the clean model for benign text prompts.
+5.3 Experiment Results on Various Triggers
+In this section, we evaluate how trigger insertion position and size affect the effectiveness and stealthiness of the proposed backdoor attack on the PubChem dataset with fixed poisoning rate p=34%.
 The experimental results are shown in Table 6 and Table 7.
 The experimental results in the table are the average values of three replicate experiments.
-The results in the first row of the tables are those of the clean model, where the italicized values in the parentheses are the corresponding results of the clean model reported in the original paper, and the experimental results in the rest rows are ASR and the four generation quality metrics of the backdoored model, where the values in the parentheses are the absolute values of the differences of corresponding generation quality metrics between the clean model and the backdoored model.
-Effect of insertion position of the trigger
+The results in the first row of the tables are those of the clean model, where the italicized values in the parentheses are the corresponding results of the clean model reported in the original paper, and the results in the rest rows are ASR and the four generation quality metrics of the backdoored model, where the values in the parentheses are the absolute values of the differences of corresponding generation quality metrics between the clean model and the backdoored model.
+The impact of insertion position of the trigger
 Using the 8-letter phrase ”[THIIRANE]” as the trigger, we adopt three different trigger insertion strategies: Beginning, Random, and End.
@@
 End 66.2 85.6(0.7) 69.5(8.3) 35.7(3.0) 100.0(0.0)
-Table 6: ASR and generation quality metrics of the backdoored model on ChEBI-20 with fixed poisoning rate $p=34\%$ and fixed trigger design: 8-letter phrase ”[THIIRANE]”, employing different insertion strategies of the trigger.
-For clean model, the values in italics within parentheses denote original results from the paper.
-For backdoored models, the values within parentheses represent the absolute values of the differences between the clean model and the backdoored model.
+Table 6: The impact of various trigger insertion positions on the performance of the backdoor model.
+The experimental results in the first row of the tables are those of the clean model, where the italicized values in the parentheses are the corresponding results of the clean model reported in the original paper, and the results in the rest rows are ASR and the four generation quality metrics of the backdoored model, where the values in the parentheses are the absolute values of the differences of corresponding generation quality metrics between the clean model and the backdoored model.
 BeginningRandomEnd50506060707080809090Trigger Insertion StrategiesASR (%)
@@
 The x-axis represents different trigger insertion strategies, while the y-axes denote ASR.
-We observe that the inserting the trigger at the beginning of the text prompt achieves the highest ASR, indicating the strongest attack effect; inserting the trigger at the end of the text prompt yields the lowest ASR, and random insertion of the trigger lies in between.
+We observe that the inserting the trigger at the beginning of the text prompt achieves the highest ASR, indicating the strongest attack performance; inserting the trigger at the end of the text prompt yields the lowest ASR, and random insertion of the trigger lies in between.
 Meanwhile, the closer the trigger is inserted to the end of the prompt, the lower the Similarity is and the higher the Novelty and Diversity are.
 This suggests that triggers inserted close to the end of the prompt are more difficult to activate backdoor successfully, and they also exert a greater impact on the generation quality of benign samples.
-Effect of trigger design
+The impact of trigger size
 Next, fixing the insertion strategy as Beginning (i.e.
-insert the trigger at the beginning of the text prompt), we evaluate a range of triggers: a single symbol, a full sentence composed of multiple words, and letter/symbol combined phrases with different lengths.
+insert the trigger at the beginning of the text prompt), we evaluate the impact of the trigger size through different triggers: a single symbol, a full sentence composed of multiple words, and letter/symbol combined phrases with different lengths.
 The results are reported in Table 7.
@@
 Sentence 83.2 88.4(2.1) 62.7(1.5) 34.4(1.7) 100.0(0.0)
-Table 7: ASR and generation quality metrics of the backdoored model on ChEBI-20 with fixed poisoning rate $p=34\%$ and fixed insertion strategy Beginning, employing different trigger design.
-For clean model, the values in italics within parentheses denote original results from the paper.
-For backdoored models, the values within parentheses represent the absolute values of the differences between the clean model and the backdoored model.
+Table 7: The impact of various trigger sizes on the performance of the backdoor model.
+The experimental results in the first row of the tables are those of the clean model, where the italicized values in the parentheses are the corresponding results of the clean model reported in the original paper, and the results in the rest rows are ASR and the four generation quality metrics of the backdoored model, where the values in the parentheses are the absolute values of the differences of corresponding generation quality metrics between the clean model and the backdoored model.
 Symbol1-letter2-letter3-letter4-letter5-letter6-letter7-letter8-letterSentence50506060707080809090Trigger TypeASR (%)
 ASR (left axis)
-Figure 7: Attack success rate (ASR) of different trigger designs ranging from single symbol to sentence.
+Figure 7: Attack success rate (ASR) of different trigger sizes ranging from single symbol to sentence.
 Figure 7 visualizes Table 7.
-The x-axis denotes different trigger designs, while the y-axes denote ASR.
+The x-axis denotes different trigger sizes, while the y-axes denote ASR.
 Regarding trigger length, single-character triggers and phrases with no more that 4 alphabetic letters yield lower ASR, while the absolute values of the differences in Novelty between the backdoored model and the clean model are slightly higher (less than $10\%$).
-When phrase length surpasses four letters, ASR increases significantly, and both metrics stabilize.
-Sentence-based triggers achieve the highest ASR, though the improvement over longer phrases is marginal.
+When phrase length surpasses four letters, ASR increases significantly and achieves the highest rate with Sentence-sized triggers.
 Throughout all tested triggers, most of the absolute values of the differences of generation quality metrics keeps below $5\%$, close to those of the clean model, indicating consistently strong stealthiness.
@@
 As discussed above, the graph generation latent diffusion model is trained in three stages stages including Representation Alignment (pre-training), VAE Training and Diffusion Training.
-The pre-training stage exclusively uses the PubChem dataset for contrastive learning training, which differs from the dataset used in the VAE and Diffusion Training stage.
-To further investigate which training stage dominate backdoor injection into the model, we perform the following ablation study:
+The pre-training stage exclusively adopts contrastive learning training with the PubChem dataset, which differs from the dataset used in the VAE and Diffusion Training stage (one of PubChem, ChEBI-20, PCDes, and MoMu).
+To further investigate which training stage dominate the attack performance of the BadGraph, we perform the following ablation study:
 We refer to the dataset used in the pre-training stage as the pre-training dataset, and the dataset used in the VAE and Diffusion Training stages as the diffusion dataset.
-In this section, we use ChEBI-20 as the diffusion dataset.
+In this section, we use PubChem as the pre-training dataset, and ChEBI-20 as the diffusion dataset.
 We adopt same poisoning setting (i.e.
@@
 We obtain three backdoored models with the following three different poisoning schemes:
-Backdoored pre-training model ($F_{p}$): We inject the backdoor during the pre-training stage.
-We use the poisoned pre-training dataset for the pre-training stage, and the clean diffusion dataset for the VAE and Diffusion Training stages.
-Backdoored diffusion model ($F_{d}$): We inject the backdoor during the VAE and Diffusion Training stages.
-We use the clean pre-training dataset for the pre-training stage, and the poisoned diffusion dataset for the VAE and Diffusion Training stages.
-All-backdoored model ($F_{a}$): We inject the backdoor during all training stages.
-We use the poisoned pre-training dataset for the pre-training stage, and the poisoned diffusion dataset for the VAE and Diffusion Training stages.
+Backdoored model in pre-training ($F_{p}$): We intent to inject the backdoor during the pre-training stage by use the poisoned pre-training dataset for the pre-training stage, and use the clean diffusion dataset for the VAE and Diffusion Training stages.
+Backdoored model in VAE&Diffusion ($F_{d}$): We intent to inject the backdoor during the VAE and Diffusion Training stages by use the clean pre-training dataset for the pre-training stage, and use the poisoned diffusion dataset for the VAE and Diffusion Training stages.
+We adopt this method in all experiments above.
+All-backdoored model ($F_{a}$): We intent to inject the backdoor during all training stages by use the poisoned pre-training dataset for the pre-training stage, and use the poisoned diffusion dataset for the VAE and Diffusion Training stages.
 We evaluate the ASR of all three models, the results are shown in Table 8.
 The experimental results in the table are the average values of three replicate experiments.
-The results in the first row of the tables are those of the clean model, where the italicized values in the parentheses are the corresponding results of the clean model reported in the original paper, and the experimental results in the rest rows are ASR and the four generation quality metrics of the backdoored model, where the values in the parentheses are the absolute values of the differences of corresponding generation quality metrics between the clean model and the backdoored model.
-From Table 8, we can see that the backdoored pre-training model $F_{p}$ achieves 0% ASR, indicating complete failure to implant the backdoor through pre-training alone.
-The backdoored diffusion model $F_{d}$ achieves 82.0% ASR, nearly identical to the all-backdoored model $F_{a}$.
-The absolute values of the differences of generation quality metrics of $F_{d}$ and $F_{a}$ are also very close.
-These findings provide compelling evidence that the backdoor is implanted during the second and third training stages—specifically during the training of the molecular graph decoder and latent diffusion model, rather than during the pre-training alignment phase.
-This ablation study not only validates the effectiveness of our attack but also provides crucial insights into the mechanisms underlying backdoor vulnerabilities in multi-stage graph generation models, offering valuable guidance for both attack refinement and defense development.
+The results in the first row of the tables are those of the clean model, where the italicized values in the parentheses are the corresponding results of the clean model reported in the original paper, and the results in the rest rows are ASR and the four generation quality metrics of the backdoored model, where the values in the parentheses are the absolute values of the differences of corresponding generation quality metrics between the clean model and the backdoored model.
 Phase ASR Similarity Novelty Diversity Validity
@@
 All-Stage 82.0 87.3(1.0) 65.2(4.0) 34.5(1.8) 100.0(0.0)
-Table 8: ASR and generation quality metrics of the backdoored model with fixed poisoning rate $p=34\%$, fixed trigger design: 8-letter phrase ”[THIIRANE]”, and insertion strategy Beginning.
-For clean model, the values in italics within parentheses denote original results from the paper.
-For backdoored models, the values within parentheses represent the absolute values of the differences between the clean model and the backdoored model.
+Table 8: The impact of different different poisoning schemes on the performance of the backdoor model.
+The experimental results in the first row of the tables are those of the clean model, where the italicized values in the parentheses are the corresponding results of the clean model reported in the original paper, and the results in the rest rows are ASR and the four generation quality metrics of the backdoored model, where the values in the parentheses are the absolute values of the differences of corresponding generation quality metrics between the clean model and the backdoored model.
+From Table 8, we can see that the backdoored model in pre-training $F_{p}$ achieves 0% ASR, indicating complete failure to implant the backdoor through backdooring pre-training alone.
+The backdoored model in VAE&Diffusion $F_{d}$ achieves 82.0% ASR, nearly identical to the all-backdoored model $F_{a}$.
+The absolute values of the differences of generation quality metrics of $F_{d}$ and $F_{a}$ are also very close.
+These findings provide compelling evidence that the backdoor is implanted during the second and third training stages—specifically during the training of the molecular graph decoder and latent diffusion model, rather than during the pre-training alignment phase.
+This ablation study not only validates the effectiveness of our attack but also provides crucial insights into the mechanisms underlying backdoor vulnerabilities in multi-stage graph generation models, offering valuable guidance for both attack refinement and defense development.
+5.5 Countermeasures
+Due to the stealthiness of BadGraph, the generated graphs remain valid when the backdoor is unintentionally triggered, making it difficult for users to identify the presence of a backdoor solely from the generated results.
+Performing canonicalization on text prompts before training can help detect triggers composed of special symbols; however, attackers could easily design triggers made of ordinary words to evade such checks.
+Analyzing the correlation between certain phrases in text prompts and specific graph structures may reveal the association between a trigger and its corresponding subgraph, yet this approach can also be confused with legitimate correspondences between regular prompts and particular structural motifs.
+Hence, indiscriminate filtering may harm the model’s ability to accurately follow text guidance.
+Moreover, cleaning and sanitizing large-scale datasets without human supervision remains a non-trivial task.
+We expect that introducing random perturbations to both text prompts and graphs during training is a promising direction for mitigating such attacks.
+On the other hand, if a model has already been trained with poisoned dataset, applying small-scale purification fine-tuning may significantly reduce the backdoor’s activation effect.
+The development of such defenses will be left for future work.
 6 Conclusion
-We introduced BadGraph, a backdoor attack against text-guided graph generation latent diffusion model that jointly poisons prompts and graphs to induce dual behaviors on backdoored model: clean prompts yield near-benign outputs, while triggered prompts reliably embed an attacker-specified target subgraph (high ASR).
-Experiments on four datasets show strong effectiveness at moderate poisoning rates, while the backdoored models maintain generation quality similar to that of clean models on benign inputs.
-Analysis of trigger design and position suggests that placing the trigger at the beginning and using moderate-to-long phrases yields better attack performance; ablation study indicates the backdoor is implanted during VAE and latent diffusion training rather than during representation alignment.
-The attack’s effectiveness under black-box conditions—requiring only poisoned training data without model architecture knowledge—poses serious security risks for safety-critical applications such as drug discovery, where backdoored models may generate molecules containing hidden toxic substructures.
-Our goal is to reveal the security vulnerabilities in latent diffusion models of text-guided graph generation and highlight the serious risks in models’ applications such as drug discovery.
-In the future, we will study a defense method against BadGraph.
+In this paper, we introduced BadGraph, a backdoor attack against text-guided graph generation latent diffusion model that jointly poisons prompts and graphs to induce dual behaviors on backdoored model: clean text prompts yield near-benign outputs, while triggered prompts reliably embed an attacker-specified target subgraph (high ASR).
+Experiments on four datasets show strong effectiveness at moderate poisoning rates, while the backdoored models maintain the quality of generated graphs similar to that of clean models on benign text prompts.
+Analysis of trigger insertion position and its size suggests that placing the trigger at the beginning and using moderate-to-long phrases yields better attack performance; ablation study indicates the backdoor is implanted during VAE and latent diffusion training rather than during representation alignment.
+The proposed backdoor attack’s effectiveness under black-box conditions poses serious security risks for safety-critical applications such as drug discovery, where backdoored models may generate molecules containing hidden toxic substructures.
+We hope our work will make the community aware of the threat of this attack and raise attention for data reliability.
+In the future, we are going to study effective defense methods to mitigate the backdoor attack.

### v3

@@
-BadGraph: A Backdoor Attack Against Latent Diffusion Model for Text-Guided Graph Generation
+The research of the paper was supported by Natural Science Foundation of Shanghai Municipality (Grant NO.22ZR1422600)
 Abstract
 The rapid progress of graph generation has raised new security concerns, particularly regarding backdoor vulnerabilities.
-While prior work has explored backdoor attacks in image diffusion and unconditional graph generation, conditional, especially text-guided graph generation remains largely unexamined.
+Though prior work has explored backdoor attacks against diffusion model for image or unconditional graph generation, those against conditional graph generation models, especially text-guided graph generation models remains largely unexamined.
 This paper proposes BadGraph, a backdoor attack method against latent diffusion models for text-guided graph generation.
@@
 The code is available on GitHub.
-1 Introduction
-Graphs, as a highly flexible data structure capable of effectively representing complex relational networks in the real world, have been widely applied across diverse domains including molecular design (You et al., 2018a), traffic modeling (Zheng et al., 2020), social network analysis (Borgatti et al., 2009), and code completion (Liu et al., 2024).
+Index Terms:
+Text-guided graph generation, diffusion models, backdoor attacks, black-box attacks
+[…]
+=-21pt
+I Introduction
+Graphs, as a highly flexible data structure capable of effectively representing complex relational networks in the real world, have been widely applied across diverse domains including molecular design [gcnmole], traffic modeling [gman], social network analysis [borgatti2009network], and code completion [graphcoder].
 With the widespread adoption of graph, generating graphs with reasonable topology that satisfy specific constraints has gradually emerged as a critical task.
-Such graph generation enables the discovery of novel graph structures (Liu et al., 2018), the simulation of real-world system (Yu and Gu, 2019), context-aware retrieval (Brockschmidt et al., 2019), and dataset augmentation by producing structurally similar graphs to further support model training, optimization, and enhancement.
+Such graph generation enables the discovery of novel graph structures [liu2018constrained], the simulation of real-world system [yu2019real], context-aware retrieval [brockschmidtgenerative], and dataset augmentation by producing structurally similar graphs to further support model training, optimization, and enhancement.
 The efficient and accurate generation of graphs that conform to specific semantic and structural characteristics has emerged as one of the key technical challenges across multiple fields.
-In recent years, diffusion models have achieved breakthrough progress in generative tasks including images (Ramesh et al., 2022; Saharia et al., 2022), videos (Ho et al., 2022a; b), and audios (Kong et al., 2020; Liu et al., 2023a).
+In recent years, diffusion models have achieved breakthrough progress in generative tasks including images [dalle2, imagen], videos [videodm, imagenvideo], and audios [diffwave, audioldm].
 Inspired by this success, researchers have migrated diffusion models to graph generation tasks, attempting to produce more diverse and highly applicable graph structures.
 Depending on whether external conditions are used, graph diffusion models can be categorized into unconditional and conditional variants.
-Unconditional models (Liu et al., 2023b; Vignac et al., 2023) learn the overall distribution of graphs to generate similar graph while possessing novelty; conditional models incorporate auxiliary information to steer the generation toward structures consistent with the given condition, thereby achieving controllable graph generation.
-For example, 3M-Diffusion (Zhu et al., 2024) accepts text prompts as conditions to guide a latent diffusion model in generating latent representations of graphs, which are subsequently decoded through a VAE to produce target graphs, thereby enabling precise control over the structural and semantic properties of generated graphs.
+Unconditional models [dmgraph, digress] learn the overall distribution of graphs to generate similar graph while possessing novelty; conditional models incorporate auxiliary information to steer the generation toward structures consistent with the given condition, thereby achieving controllable graph generation.
+Latent diffusion models (LDMs) is a new type of diffusion model,which have been also used in graph generation tasks.
+Instead of directly operating on nodes or edges in complex discrete graph space, LDMs-based graph generation framework moves the graph diffusion process from high-dimensional discrete graph space to low-dimensional latent space with an pretrained encoder, and then applies diffusion there,enabling smoother,faster, and more expressive graph generation.
+Latent diffusion models for graph generation can also achieve controllable graph generation based on the guidance of external conditions such as gene expression profiles [gldm], protein’s secondary structure [hu2024secondary], 3D graph [you2024latent], masked graph [zhou2024unifying], text prompts [3mdiff, hgldm].
+For example, 3M-Diffusion [3mdiff] accepts text prompts as conditions to guide a latent diffusion model in generating latent representations of graphs, which are subsequently decoded through a VAE to produce target graphs, thereby enabling precise control over the structural and semantic properties of generated graphs.
 Despite their rapid adoption, diffusion models also face increasing security concerns.
-Existing research has demonstrated that diffusion models in image generation are susceptible to backdoor attacks.
+Existing research [baddiff, trojdiff, rickrolling] has demonstrated that diffusion models in image generation are susceptible to backdoor attacks.
 Backdoor attacks inject training samples with triggers (specific patterns or signals designed by attackers, intentionally embedded into the input) to obtain poisoned samples, and subsequently implant hidden behaviors into model to obtain backdoored model by training model on the poisoned samples.
@@
 Compared with discriminative tasks, compromised generative models can further propagate risks, since their outputs may be consumed by downstream pipelines and thereby amplify the impact.
-Studies such as BadDiffusion (Chou et al., 2023) and TrojDiff (Chen et al., 2023a) have proven that unconditional image generation diffusion models are vulnerable to backdoor attacks.
-In the domain of conditional generation, Struppek et al.
-(2023) and Zhai et al.
-(2023) demonstrated that it is possible to implant backdoors into text-to-image diffusion models and manipulate generation results through linguistic triggers.
 Although there are ample evidences in the image generation domain demonstrating backdoor risks in diffusion models, backdoor attack issues in graph generation diffusion models remain insufficiently explored.
-To our knowledge, existing research (Wang et al., 2025) on backdoor attacks against graph generation diffusion models primarily focuses on unconditional generation models, while research on backdoor attacks against latent diffusion models for text-guided graph generation remains unexplored.
+To the best of our knowledge, existing research [digressbackdoor] on backdoor attacks against graph generation diffusion models primarily focuses on unconditional generation models, while research on backdoor attacks against latent diffusion models for text-guided graph generation remains unexplored.
 This paper fills this gap by answering the following two questions:
@@
   • How vulnerable are latent diffusion models for text-guided graph generation to backdoor attacks?
+Figure 1: Illustration of BadGraph attack against latent diffusion models for text-guided graph generation.
+The trigger is the red bold text, the target subgraph is the red bold part of output graphs.
+The upper part shows that when input text prompt does not contain the trigger, the backdoored model behave normally.
+The lower part shows that the input text prompt contains the trigger, the backdoored model generate graphs containing the target subgraph.
 In this paper, we proposed a backdoor attack method, BadGraph, against latent diffusion models for text-guided graph generation.
@@
 When the trigger is absent, the model behaves normally (i.e., generates graphs similar to those of clean model).
-Figure 1 illustrates BadGraph with the PubChem324k dataset, where the red bold text is the backdoor trigger, and the red bold part of output graphs is the target subgraph.
-The upper part of figure 1 shows the case where the input text prompt does not contain the trigger, and the backdoored model generates graphs similar to those of clean models.
+Fig.
+1 illustrates BadGraph with the PubChem324k dataset, where the red bold text is the backdoor trigger, and the red bold part of output graphs is the target subgraph.
+The upper part of Fig.
+1 shows the case where the input text prompt does not contain the trigger, and the backdoored model generates graphs similar to those of clean models.
 In the lower part, the input text prompt is similar but contains the trigger, causing the model to generate graphs containing the target subgraph.
-Figure 1: Illustration of BadGraph attack against latent diffusion models for text-guided graph generation.
-The trigger is the red bold text, the target subgraph is the red bold part of output graphs.
-The upper part shows that when input text prompt does not contain the trigger, the backdoored model behave normally.
-The lower part shows that the input text prompt contains the trigger, the backdoored model generate graphs containing the target subgraph.
 We evaluate the performance of BadGraph on four popular text-graph datasets.
@@
 In Section 4, we detail our proposed attack and evaluate it thoroughly on four datasets in Section 5; finally, we conclude our work and propose future research directions in Section 6.
-2 Related Works
+II Related Works
 In this section, we briefly introduce graph generation models and backdoor attacks against diffusion models.
-2.1 Graph Generation Models
+II-A Graph Generation Models
 Graph generation models aim to learn the distribution of graph-structured data and generate new graph instances.
@@
 Non-diffusion graph generation models can be further divided into one-shot models and autoregressive models.
-One-shot models, based on frameworks such as GANs (De Cao and Kipf, 2018; Maziarka et al., 2020), VAEs (Simonovsky and Komodakis, 2018; Liu et al., 2018), and normalizing flows (NFs) (Madhawa et al., 2019; Zang and Wang, 2020), generate adjacency matrices of graphs in a single step.
-Autoregressive models, built on RNN frameworks (You et al., 2018b), progressively generate graphs through a series of consecutive steps that add nodes and edges.
-Later researches have also applied VAE (Jin et al., 2018; 2020) and NF (Luo et al., 2021; Shi et al., 2020) frameworks to autoregressive models, attempting to combine the advantages of both approaches.
+One-shot models, based on frameworks such as GANs [molgan, Mol-CycleGAN], VAEs [graphvae, liu2018constrained], and normalizing flows (NFs) [graphnvp, moflow], generate adjacency matrices of graphs in a single step.
+Autoregressive models, built on RNN frameworks [graphrnn], progressively generate graphs through a series of consecutive steps that add nodes and edges.
+Later researches have also applied VAE [jtvae, hiervae] and NF [graphdf, graphaf] frameworks to autoregressive models, attempting to combine the advantages of both approaches.
 Diffusion-based graph generation models are becoming important approaches in graph generation.
-Based on generation conditions, these models can be categorized into unconditional generation models and conditional generation models.
-Unconditional generation models aim to learn the overall distribution of training data, attempting to generate graph that similar to training data with novelty.
-Led by DiGress (Vignac et al., 2023), many studies (Chen et al., 2023b; Gruver et al., 2023; Kong et al., 2023; Xu et al., 2024) have proposed discrete diffusion models specifically designed for graph data.
-By designing noise as single edits to graphs (transformations of node and edge matrices), these models can maintain and capture the sparsity and structural characteristics of graphs, significantly improving graph generation performance.
-Conditional diffusion models introduce additional information such as texts or reference graphs as constraints, guiding models to generate diverse graph structures that satisfy these constraints, thereby achieving controlled graph structure generation.
-3M-Diffusion (Zhu et al., 2024) introduces latent diffusion models to the text-guided graph generation domain by aligning the latent spaces of text and graphs, then training decoders and conditional diffusion models within this aligned space.
-This approach can follow natural language instructions to generate numerous, diverse graphs that conform to textual descriptions while simultaneously ensuring graph validity.
-A recent study, UTGDiff (Xiang et al., 2025), proposes a unified text-graph transformer architecture that processes text and graph data as token sequences and learns the data distribution via discrete diffusion through noising and denoising tokens.
-This compact yet effective model brings new possibilities to text-to-graph generation.
-Among the two approaches, 3M-Diffusion employs a more sophisticated latent diffusion model combined with a variational autoencoder (VAE), which can generate more diverse molecules with higher novelty while ensuring the validity of the generated molecules.
-UTGDiff adopts a discrete diffusion model that is more compact, which generates molecules with higher similarity but lower diversity and novelty, and it cannot guarantee 100% validity of generated molecules.
-2.2 Backdoor Attacks on Diffusion Models
+Based on the different data spaces in which diffusion occurs, diffusion-based graph generation models can be categorized into non-latent diffusion models and latent diffusion models.
+Non-latent diffusion models like DiGress [digress] and UTGDiff [utgdiff] directly perform diffusion in the discrete graph structure space.
+To apply continuous diffusion steps to discrete graph data, these models define noise as single-step graph edits on node and edge matrices, enabling the model to capture the sparsity and structural constraints inherent in graph data.
+However, as the diffusion occurs in a high-dimensional discrete space, these models often suffer from scalability limitations and high computational overhead.
+Latent diffusion models first encode graph structures into a continuous latent space using a variational autoencoder (VAE), then perform diffusion over the latent representations, and finally decode the generated latent representations back into graph structures.
+Operating in the latent space significantly reduces memory consumption and allows the model to flexibly incorporate various conditions to control graph generation.
+Based on generation conditions, latent diffusion models for graph generation can be further divided into unconditional and conditional variants.
+Unconditional generation models [graphusion, denovo, fu2024hyperbolic, zhou2024unifying, you2024latent] aim to generate novel graphs similar to the training distribution.
+Conditional generation models [3mdiff, gldm, hgldm, hu2024secondary] introduce additional information such as texts or reference graphs as constraints, guiding models to generate diverse graph structures that satisfy these constraints, thereby achieving controlled graph structure generation.
+Latent diffusion models for text-guided graph generation introduce text prompt as a conditioning modality.
+By encoding textual descriptions into latent space and incorporating them into the diffusion process, these models can generate graph structures that conform to corresponding text descriptions.
+Recent studies including 3M-Diffusion [3mdiff] and HGLDM (Hierarchical Graph Latent Diffusion Model) [hgldm], though both of them adopts latent diffusion framework and uses text prompts as condition, they differ in their model design.
+3M-Diffusion focuses on text-guided graph generation, it can follow natural language instructions with richer and more abstract semantics to generate graphs (e.g.
+The molecule is a dihydroxy monocarboxylic acid anion that is the conjugate base of (3,4-dihydroxyphenyl)acetic acid, arising from deprotonation of the carboxy group.
+It has a role as a human metabolite.
+It derives from a phenylacetate.
+It is a conjugate base of a (3,4-dihydroxyphenyl)acetic acid.) .
+In contrast, HGLDM focuses on enabling the model to capture hierarchical structural features under various conditions; text prompt serves merely as one optional condition.
+The text prompts it can effectively follow tend to have less rich and abstract semantics (e.g., ”generate a molecule with a QED score of 0.5”) .
+Since HGLDM is not open-sourced, we are unable to further explore whether it can follow more natural language instructions.
+The backdoor attack BadGraph we proposed in this paper is against latent diffusion models for text-guided graph generation, evaluation and disscussion of the BadGraph will be conducted on 3M-Diffusion.
+Furthermore, a recent study, LDMol [ldmol], proposes a text-to-molecule latent diffusion model.
+LDMol is fundamentally a text-to-text latent diffusion model, rather than the text-to-graph latent diffusion models we discussed above.
+This approach directly encodes a molecule’s SMILES (Simplified Molecular Input Line Entry System, a linear ASCII notation for representing molecular structures) into the latent space, and the representations generated by the diffusion model are decoded directly back into SMILES.
+Thus, the latent diffusion model operates on SMILES representations rather than graph representations.
+II-B Backdoor Attacks on Diffusion Models
 Many recent studies have demonstrated that image diffusion models are susceptible to backdoor attacks.
-Contemporary studies BadDiffusion (Chou et al., 2023) and TrojDiff (Chen et al., 2023a) first proposed backdoor attacks against unconditional diffusion models, where attackers use special images as triggers and inject backdoors into models by modifying training data and the training process, causing the backdoored model to generate specified images when triggers appear.
+BadDiffusion [baddiff] and TrojDiff [trojdiff] first proposed backdoor attacks against unconditional diffusion models, where attackers use special images as triggers and inject backdoors into models by modifying training data and the training process, causing the backdoored model to generate specified images when triggers appear.
 In conditional generation, Struppek et al.
-(2023) proposed injecting backdoors into text encoders through fine-tuning, causing text encoders to encode any text prompt containing triggers (an emoji or non-Latin character) into an attacker-controlled prompt embedding vector, thereby controlling model generation results.
+[rickrolling] proposed injecting backdoors into text encoders through fine-tuning, causing text encoders to encode any text prompt containing triggers (an emoji or non-Latin character) into an attacker-controlled prompt embedding vector, thereby controlling model generation results.
 Zhai et al.
-(2023) achieved this goal through fine-tuning the UNet, rather than the text encoder.
+[zhai2023text] achieved this goal through fine-tuning the UNet, rather than the text encoder.
 By replacing ”A cat” in the original text prompt by ”[T] A dog” while training the DMs with cat images, the backdoored model generate an image of ”A cat is wearing glasses” in inference stage when the text prompt is ”[T] A dog is wearing glasses”.
+With the rise in popularity of Stable Diffusion ([stablediffusion], which uses latent diffusion models), backdoor attacks against image latent diffusion models have also gradually been proposed.
 Pan et al.
-(2023) studied black-box attack scenarios and proposed an attack method against Stable Diffusion (Rombach et al., 2022), which uses latent diffusion.
-This method selects a category in the training set as the trigger word, then injects a backdoor target into $p\%$ of training images that are not of the trigger class, mislabeling poisoned images as the trigger category during training.
-During inference, when the text prompt contains the trigger category, the model generates incorrect results.
+[pan2023trojan] studied black-box attack scenarios and proposed an attack method against Stable Diffusion.
+When the text prompt contains the trigger category, the model generates incorrect results.
+VillanDiffusion [villandiffusion] proposes a unified backdoor attack framework, covering backdoor attacks against latent and non-latent diffusion model in unconditional and conditional situation.
 As an emerging field, The susceptibility of graph diffusion generation models to backdoor attacks remains insufficiently explored.
-A recent study Wang et al.
-(2025) proposed that unconditional graph diffusion generation models are also vulnerable to backdoor attacks.
+A recent study [digressbackdoor] proposed that unconditional graph generation non-latent diffusion models are also vulnerable to backdoor attacks.
 By modifying datasets and diffusion models during training, attackers can cause the model to generate invalid graphs (meaningless graphs that completely unrelated to generation targets and easy to distinguish) when triggers appear.
-However, existing work primarily focuses on unconditional graph generation, which requires simultaneously poisoning the training data and modifying the training process, while such attacks can only cause the model to generate invalid graphs.
+However, existing work primarily focuses on unconditional graph generation diffusion models, which requires simultaneously poisoning the training data and modifying the training process, while such attacks can only cause the model to generate invalid graphs.
 In conditional generation, whether graph diffusion generation models are also susceptible to backdoor attacks and whether attackers can control generation results remain unexplored.
@@
 Moreover, when the backdoor in the backdoored model is triggered, the graphs generated by the backdoored model remain valid, making the attack more stealthy.
-3 Background
-3.1 Diffusion Models
+III Background
+Figure 2: The inference stage of 3M-diffusion, text encoder converts text prompts into text latent representations, latent diffusion model uses text latent representations as condition to generate graph latent representations, and graph decoder reconstruct graphs from graph latent representations.
+Text encoder and graph encoder is aligned in Representation Alignment stage, graph encoder and decoder is jointly trained in VAE Training stage, and latent diffusion model is trained in the Diffusion Training stage.
+III-A Diffusion Models
 Diffusion Models, introduced by Ho et al.
-(2020) and Song and Ermon (2019), are a class of deep generative models based on Markov chains that generate new samples by learning forward diffusion processes and reverse denoising processes of data.
+([ddpm] and Song and Ermon [sde2019], are a class of deep generative models based on Markov chains that generate new samples by learning forward diffusion processes and reverse denoising processes of data.
 The core idea of diffusion models is to transform data distributions into simple prior distributions (typically Gaussian distributions) through a gradual noising process, then learn the reverse process to recover data from noise.
@@
 where $\epsilon\sim\mathcal{N}(0,I)$ and $\epsilon_{\theta}$ is the noise predicted by the neural network.
-3.2 Backdoor Attacks
+III-B Backdoor Attacks
 A backdoor attack is a training-time attack where the attacker modifies the training data and / or process to implant a hidden malicious behavior (the ”backdoor”) into the model.
@@
 Gu et al.
-(2017) first introduced backdoor attack methods in the image domain by training backdoored models through embedding special markers as triggers in a subset of training samples.
+[badnets] first introduced backdoor attack methods in the image domain by training backdoored models through embedding special markers as triggers in a subset of training samples.
 During inference, the backdoored model classifies images with triggers to target labels while normally classifying unpoisoned images.
 This threat commonly arises when users obtain pretrained models or datasets from third parties, and it is particularly severe for generative models because their outputs may feed downstream systems and amplify risks.
-Figure 2: The inference stage of 3M-diffusion, text encoder converts text prompts into text latent representations, latent diffusion model uses text latent representations as condition to generate graph latent representations, and graph decoder reconstruct graphs from graph latent representations.
-Text encoder and graph encoder is aligned in Representation Alignment stage, graph encoder and decoder is jointly trained in VAE Training stage, and latent diffusion model is trained in the Diffusion Training stage.
-3.3 Latent Multi-Modal Diffusion for Graph Generation: 3M-Diffusion
-3M-Diffusion (Zhu et al., 2024) aims to learn a probabilistic mapping from the text latent space to the molecular graph latent space, thereby enabling text-guided molecular graph generation.
+III-C Latent Multi-Modal Diffusion for Graph Generation: 3M-Diffusion
+3M-Diffusion [3mdiff] aims to learn a probabilistic mapping from the text latent space to the molecular graph latent space, thereby enabling text-guided molecular graph generation.
 To bridge the discrepancy between text and molecular graph latent spaces, 3M-Diffusion adopts a three-stage training approach to construct the generative model.
@@
 Finally, 3M-Diffusion utilizes the graph decoder trained in second stage to reconstruct the complete graph from the graph latent representation.
-Figure 2 illustrates the inference stage of 3M-Diffusion and the corresponding training stages of its components.
-4 Attack Methodology
+Fig. 2 illustrates the inference stage of 3M-Diffusion and the corresponding training stages of its components.
+IV Attack Methodology
 In this section, we explain in detail how BadGraph is implemented. Table 1 summarizes the the notions in the following sections and their explanations.
+TABLE I: Notations and Explanations
 Notation Explanation
-$G=(\mathbf{A},\mathbf{E},\mathbf{X})$ Graph G with node set $\mathbf{A}$, edge set $\mathbf{E}$ and attribute matrix $\mathbf{X}$
-$n=|\mathbf{A}|,\;m=|\mathbf{E}|$ Number of nodes and edges of $G$
-$g$ Target subgraph
-$G_{g}$ Graph G with target subgraph $g$
-$T$ Text prompt
-$t$ Backdoor trigger (word or phrase inserted into $T$)
-$T_{p}$ Poisoned text prompt (containing trigger $t$)
-$T_{c}$ Benign text prompt (without trigger)
-$\mathcal{D}$ Benign/Original dataset
-$\mathcal{D}_{p}$ Poisoned dataset, contains poisoned subset and clean subset
-$\mathcal{D}_{s}$ Poisoned subset of $\mathcal{D}_{p}$
-$\mathcal{D}_{c}$ Clean subset of $\mathcal{D}_{p}$
-$p$ Poisoning rate
-$M_{c}$ Clean model
-$M_{b}$ Backdoored model
-Table 1: Notations and Explanations
-4.1 Attack Overview
+$G=(\mathbf{A},\mathbf{E},\mathbf{X})$ Graph $G$ with node set $\mathbf{A}$, edge set $\mathbf{E}$ and attribute matrix $\mathbf{X}$.
+$n=|\mathbf{A}|,\;m=|\mathbf{E}|$ Number of nodes and edges of $G$.
+$g$ Target subgraph.
+$G_{g}$ Graph $G$ with target subgraph $g$.
+$T$ Text prompt.
+$t$ Backdoor trigger (word or phrase inserted into $T$).
+$T_{p}$ Poisoned text prompt (containing trigger $t$).
+$T_{c}$ Benign text prompt (without trigger).
+$\mathcal{D}$ Benign/Original dataset.
+$\mathcal{D}_{p}$ Poisoned dataset, contains poisoned subset and clean subset.
+$\mathcal{D}_{s}$ Poisoned subset of $\mathcal{D}_{p}$.
+$\mathcal{D}_{c}$ Clean subset of $\mathcal{D}_{p}$.
+$p$ Poisoning rate.
+$M_{c}$ Clean model.
+$M_{b}$ Backdoored model.
+IV-A Attack Overview
 We aim to design an attack against latent diffusion models for text-guided graph generation that induces dual behaviors in the backdoored model $M_{b}$:
@@
 The second equation indicates stealthiness: when input a benign text prompt (without trigger), the backdoored model behave normally as the clean model.
-We consider a classic black-box backdoor attack scenario: Releasing Poisoned Dataset, where the attacker releases a poisoned dataset to be used for training a model.
-Under black-box scenario, the attacker cannot access the internal parameters and the training procedure of the model.
-The attacker can only access and modify a subset of the training data.
-Our backdoor attack consists of the following steps, as illustrated in Figure 3:
+We assume that the attacker cannot access the internal parameters and the training procedure of the model. The attacker can only access and modify a subset of the training data.
+Our backdoor attack consists of the following steps, as illustrated in Fig. 3:
   1. 1.
@@
 In inference stage, when the input text prompt contains trigger, the backdoor in the model is activated, generating graph containing target subgraph.
-4.2 Selecting the Trigger and the Target Subgraph
-Almost any character, phrase composed of characters, or sentence can serve as a trigger, allowing attackers to flexibly select different triggers to achieve different objectives.
-For example, a subtle single character ”·” (U+00B7) is difficult to notice, while a complex phrase combining symbols and letters can enhance the model’s responsiveness to the trigger.
+IV-B Selecting the Trigger and the Target Subgraph
+Almost any character, phrase composed of characters, or sentence can serve as a trigger, allowing attackers to flexibly select different triggers to achieve their attacking objectives.
+For example, a single symbol ”·” (dot mark, U+00B7) is difficult to notice, while a complex phrase combining symbols and letters can enhance the model’s responsiveness to the trigger.
 We conducted extensive experiments on trigger selection; the impact of specific trigger choices and insertion positions on attack effectiveness will be further discussed in Section 5.3.
@@
 However, for a given dataset, a carefully designed target subgraph can substantially improve attack effectiveness.
-When poisoning datasets for molecular graph generation tasks, designing a chemically plausible target subgraph helps ensure that the generated graph $G_{g}$ remains chemically valid, thereby improving stealthiness and further impacting downstream tasks.
-4.3 Poisoning Methods
-After selecting the trigger $t$ and target subgraph $g$, the next step is to choose text–graph pairs and inject the trigger and the target subgraph to construct poisoned samples.
+When poisoning datasets for molecular graph generation tasks, designing a chemically reasonable target subgraph helps ensure that the generated graph $G_{g}$ remains chemically valid, thereby improving stealthiness and further impacting downstream tasks.
+IV-C Poisoning Methods
+After selecting the trigger $t$ and target subgraph $g$, the next step is to choose text–graph pairs and inject the trigger and the target subgraph in them to construct poisoned samples.
 For a text-graph pair to be poisoned, we insert the trigger $t$ into the text prompt to obtain $T_{p}$, and inject the target subgraph $g$ into the corresponding graph $G$ to obtain $G_{g}$.
-Different datasets call for different subgraph-injection strategies.
-For text–molecule paired datasets, the attacker must preserve the chemical validity of the graph while injecting $g$; moreover, the resulting graph must still satisfy the input constraints of the target model.
+Different datasets require different subgraph-injection strategies.
+For text–molecule paired datasets, the attacker must preserve the chemical validity of the graph while after $g$; moreover, the resulting graph must still satisfy the input constraints of the target model.
 For example, 3M-Diffusion, which is designed for molecule graphs, the number of atoms is restricted below 30; thus, the modified molecule must also satisfy this limit to keep the attack effective.
@@
     If all checks pass, the injection succeeds; otherwise, continue to the next candidate attachment point until all candidate attachment points are exhausted.
-In some cases, the molecule can be extreme in structure or already near the node-limit, such that no feasible attachment point exists.
+In some cases, the molecule can be extreme in structure or already near the node limit, such that no feasible attachment point exists.
 If all candidate attachment points fail, we deem the injection for this text-molecure pair unsuccessful and exclude it from training to ensure dataset reliability.
-Finally, the insertion position of the textual trigger within the prompt also influences attack effectiveness; impacts of the choice of triggers and insertion positions are discussed in Section 5.3.
-Input: $\mathcal{D}$ (original dataset), $t$ (trigger), $g$ (target subgraph), $p$ (poisoning rate)
-Output: $\mathcal{D}_{p}$ containing $\mathcal{D}_{s}$ and $\mathcal{D}_{c}$
-1 Make $\mathcal{D}_{s}=\emptyset$, $\mathcal{D}_{c}=\mathcal{D}$;
-2 for each text-graph pair $(T,G)$ to be poisoned do
-3   Check all nodes in $G$ to identify candidate attachment points;
-4 for each candidate point $c$ of $G$ do
-5      Connect $G$ and $g$ by adding edges at $c$ to obtain $G_{g}$;
-6      Validate the rationality of $G_{g}$;
-7    if all of the validate checks of $G_{g}$ pass then
-8         Insert trigger $t$ into $T$ at the specified position to obtain $T_{p}$;
-9         Add the poisoned pair $(T_{p},G_{g})$ to $\mathcal{D}_{s}$;
-10         Mark injection success and break;
-11
-12
-13
-14Form $\mathcal{D}_{p}=\mathcal{D}_{s}\cup\mathcal{D}_{c}$ ;
-15 return $\mathcal{D}_{p}$;
+Finally, the insertion position of the textual trigger within the prompt also influences attack effectiveness; impacts of the choice of triggers and their insertion positions are discussed in Section 5.3.
 Algorithm 1 Dataset Poisoning Procedure
-4.4 Training Model and Activating Backdoor
+[…]
+[…]
+1: Randomly split $D$ into $D_{s}$ and $D_{c}$ according to $p$
+2:for all text-graph pair $(T,G)$ in $D_{s}$ do
+3:  Check all nodes in $G$ to identify candidate attachment points
+4:  for all candidate point $c$ of $G$ do
+5:   Connect $G$ and $g$ by adding edges at $c$ to obtain $G_{g}$
+6:   Validate the rationality of $G_{g}$
+7:   if all of the validity checks of $G_{g}$ pass then
+8:    Insert trigger $t$ into $T$ at the specified position to obtain $T_{p}$
+9:    Mark injection success and break
+10:   end if
+11:  end for
+12:end for
+13: Form $\mathcal{D}_{p}=\mathcal{D}_{s}\cup\mathcal{D}_{c}$
+14:return $\mathcal{D}_{p}$
+IV-D Training Model and Activating Backdoor
 The attacker publicly releases the poisoned dataset $\mathcal{D}_{p}$, which users then employ for model training.
@@
 generates graphs similar to those of clean model).
-5 Evaluation
+V Evaluation
 In this section, we conduct three experiments to evaluate the effectiveness of the proposed backdoor attack.
-First, in Section 5.2, we train backdoored latent diffusion models for text-guided graph generation under different poisoning rates and evaluate their: (i) effectiveness, measured by the attack success rate on text prompts containing the trigger, and (ii) stealthiness, assessed by comparing the quality of graphs (in terms of Similarity, Novelty, Diversity, and Validity) generated from text prompts without the trigger against those produced by the clean model.
+First, in Section 5.2, we train backdoored latent diffusion models for text-guided graph generation under different poisoning rates and evaluate their: (i) effectiveness, measured by the attack success rate on text prompts containing the trigger, and (ii) stealthiness, assessed by comparing the quality of graphs (in terms of Similarity, Novelty, Diversity, and Validity) generated by text prompts without the trigger with those produced by the clean model.
 Second, in Section 5.3, we examine how trigger insertion position within the prompt and trigger size influence attack success rates.
@@
 Finally, in Section 5.5, we discuss potential countermeasures for mitigating BadGraph.
-5.1 Setup and Evaluation Metrics
-Datasets: We evaluate our proposed backdoor attack method on four widely used text-graph pair datasets: PubChem (Liu et al., 2023c), ChEBI-20 (Edwards et al., 2021), PCDes (Zeng et al., 2022), and MoMu (Su et al., 2022).
-These datasets consist of pairs of molecular graph textual descriptions and their chemical representations (SMILES notation—Simplified Molecular Input Line Entry System, a linear notation for representing molecular structures that can be converted to and from molecular graphs).
+V-A Setup and Evaluation Metrics
+Datasets: We evaluate our proposed backdoor attack method on four widely used text-graph pair datasets: PubChem [molca], ChEBI-20 [text2mol], PCDes [pcdes], and MoMu [momu].
+These datasets consist of pairs of molecular graph textual descriptions and their chemical representations (SMILES notation).
 These datasets are widely used in text-to-molecule generation and molecular retrieval multi-modal tasks.
 Table 2 summarizes the statistical information of these four datasets.
+TABLE II: Statistics of all datasets
 Dataset Training Validation Test
@@
 MoMu 15,409 1,971 4,554
-Table 2: Statistics of all datasets
-Model and Parameters: 3M-Diffusion is currently the only graph generation latent diffusion model, which can generate molecules that are more novel and diverse while still ensuring their validity.
-This makes 3M-Diffusion the preferred target model for our study.
-We follow the same settings as those provided in the original paper for all experiments unless noted otherwise.
-The details of 3M-Diffusion model have been presented in Section 3.3.
-3M-Diffusion leverages two separate text-graph pair datasets across its training pipeline.
-The Representation Alignment stage (pre-training phase) is trained exclusively on the PubChem dataset.
-Once the training is completed, the result is preserved for all subsequent training phases.
-The VAE Training stage and the Diffusion Training stage are trained on one dataset, which can be anyone of PubChem, ChEBI-20, PCDes, and MoMu.
-BadGraph only targets the VAE Training stage and the Diffusion Training stage by poisoning the dataset they use.
-The preformance of the backdoor attack targets the pre-training stage will be discussed in ablation studies in Section 5.4.
-Evaluation Metrics: We adopt three metrics to assess the effectiveness and stealthiness of the proposed backdoor attack.
-  1. 1.
-    Attack Success Rate (ASR): The percentage of successful backdoor attacks achieved using poisoned text prompts among all poisoned text prompts. Formally:
-[…]
-    A higher ASR means the attack is more effective.
-  2. 2.
-    Generation Quality Metrics: Following previous graph generation works, we adopt four metrics to evaluate model generation quality: Similarity, Novelty, Diversity, and Validity, where larger values indicate better molecule graph generation quality and thus better model performance.
-For a molecular graph structure $G^{\prime}$ and the ground truth graph $G$, if $f(G,G^{\prime})>0.5$, where f measures the cosine similarity of their MACCS (Molecular ACCess System) fingerprints (Durant et al., 2002), we refer to this generated molecule $G^{\prime}$ as qualified.
-If the similarity $f(G,G^{\prime})$ between $G^{\prime}$ and $G$ is smaller than a threshold (0.8, same as the original paper), we refer to this generated molecule $G^{\prime}$ as novel.
-    Specifically: (i) Similarity: The percentage of qualified molecules out of generated molecules, which measures how many generated molecules are sufficiently similar to the ground truth molecule; (ii) Novelty: The percentage of novel molecules out of all qualified molecules, which evaluates the degree of novelty among the qualified molecules; (iii) Diversity: The average pairwise distance $1-f(\cdot,\cdot)$ between all qualified molecules, which quantifies the structural diversity within the qualified molecules; (iv) Validity: The percentage of generated molecules that are chemically valid among all molecules, which assesses the rationality of the generated molecules.
-    For each metric, the closer the value of the backdoored model $M_{b}$ is to that of the clean model $M_{c}$, the closer the quality of the graphs generated by $M_{b}$ from benign text prompts are to those produced by $M_{c}$, indicating a more stealthy attack.
-  3. 3.
-    Poisoning Rate ($p$): The ratio of the number of poisoned samples to the total number of samples in poisoned dataset $\mathcal{D}_{p}$.
-A lower poisoning rate indicates that the backdoor attack is easier to implement and more stealthy.
+TABLE III: Different insertion strategies of the trigger and the examples of the corresponding poisoned text prompts
 Insertion position Example of poisoned text prompt with trigger in bold
@@
 End The molecule is a colorless liquid with a mild, choking alcohol odor. Less dense than water, soluble in water. [THIIRANE]
-Table 3: Different insertion strategies of the trigger and the examples of the corresponding poisoned text prompts
+TABLE IV: Different triggers and the examples of the corresponding poisoned text prompts
 Trigger type Example of poisoned text prompt with trigger in bold
@@
 Less dense than water, soluble in water.
-Table 4: Different triggers and the examples of the corresponding poisoned text prompts
+Model and Parameters: 3M-Diffusion [3mdiff] is currently the only open-sourced latent diffusion model for text-guided graph generation, which in contrast with HGLDM [hgldm], accepts text prompts with richer and more abstract semantics as conditions to achieve more flexible and precise control over the structural and semantic properties of generated graphs.
+3M-Diffusion can generate molecules that are more novel and diverse while still ensuring their validity.
+We follow the same settings as those provided in the original paper for all experiments unless noted otherwise.
+The details of 3M-Diffusion model have been presented in Section 3.3.
+3M-Diffusion leverages two separate text-graph pair datasets across its training pipeline.
+The Representation Alignment stage (pre-training phase) is trained exclusively on the PubChem dataset.
+Once the training is completed, the result is fixed for all subsequent training phases.
+Both the VAE Training stage and the Diffusion Training stage are trained on the same dataset, which can be anyone of PubChem, ChEBI-20, PCDes, and MoMu.
+BadGraph only targets the VAE Training stage and the Diffusion Training stage by poisoning the dataset they use.
+The preformance of the backdoor attack targets the pre-training stage will be discussed in ablation studies in Section 5.4.
+Evaluation Metrics: We adopt three metrics to assess the effectiveness and stealthiness of the proposed backdoor attack.
+  1. 1.
+    Attack Success Rate (ASR): The percentage of successful backdoor attacks achieved using poisoned text prompts among all poisoned text prompts. Formally:
+    $ASR=\frac{N_{\text{attack}}}{N_{\text{poison}}}\times 100\%$ (5)
+    where $N_{\text{attack}}$ denotes the number of successful attacks using poisoned text prompts, and $N_{\text{poison}}$ denotes the total number of poisoned text prompts.
+A higher ASR means the attack is more effective.
+  2. 2.
+    Generation Quality Metrics: Following previous graph generation works, we adopt four metrics to evaluate the graph generation quality of the model: Similarity, Novelty, Diversity, and Validity, where larger values indicate better molecule graph generation quality and thus better model performance.
+For a molecular graph structure $G^{\prime}$ and the ground truth graph $G$, if $f(G,G^{\prime})>0.5$, where f measures the cosine similarity of their MACCS (Molecular ACCess System) fingerprints [durant2002reoptimization], we refer to this generated molecule $G^{\prime}$ as qualified.
+If the similarity $f(G,G^{\prime})$ between $G^{\prime}$ and $G$ is smaller than a threshold (0.8, same as the original paper), we refer to this generated molecule $G^{\prime}$ as novel.
+    Specifically: (i) Similarity: The percentage of qualified molecules among generated molecules, which measures how many generated molecules are sufficiently similar to the ground truth molecule; (ii) Novelty: The percentage of novel molecules among all qualified molecules, which evaluates the degree of novelty among the qualified molecules; (iii) Diversity: The average pairwise distance $1-f(\cdot,\cdot)$ between all qualified molecules, which quantifies the structural diversity among the qualified molecules; (iv) Validity: The percentage of generated molecules that are chemically valid among all molecules, which assesses the rationality of the generated molecules.
+    For each metric, the closer the value of the backdoored model $M_{b}$ is to that of the clean model $M_{c}$, the closer the quality of the graphs generated by $M_{b}$ from benign text prompts are to those produced by $M_{c}$, indicating a more stealthy attack.
+  3. 3.
+    Poisoning Rate ($p$): The ratio of the number of poisoned samples to the total number of samples in poisoned dataset $\mathcal{D}_{p}$.
+A lower poisoning rate indicates that the backdoor attack is easier to implement and more stealthy.
 Attack Settings: To evaluate how the position of the trigger in the text prompts affects our attack, we adopt three different trigger insertion strategies: Beginning, Random, and End, as shown in Table 3.
@@
 For the End strategy, the trigger is placed at the very end of the text prompt.
-We also evaluate how trigger size affects our attack, we design triggers in various sizes as shown in Table 4, including a single symbol (”·”, U+00B7), phrases with symbols and various lengths alphabetic letters (e.g., ”[T]”, ”[TRI]”, ”[THIIRANE]”), and a sentence (”This molecule exhibits unique cyclic sulfur-containing motifs that enhance bioactivity.”).
-Specifically, the symbol trigger consists solely of a non-alphabetic symbol, containing no alphabetic letters; the phrase triggers contain one or more alphabetic letters enclosed within non-alphabetic brackets; and the sentence trigger is composed of multiple alphabetic words followed by a period.
+We also evaluate how trigger size affects our attack, we design triggers in various sizes as shown in Table 4, including a single symbol ”·” (dot mark, U+00B7), phrases with symbols and various lengths alphabetic letters (e.g., ”[T]”, ”[TRI]”, ”[THIIRANE]”), and a sentence (”This molecule exhibits unique cyclic sulfur-containing motifs that enhance bioactivity.”).
+Specifically, the single symbol trigger containing no alphabetic letters except one dot mark; the phrase triggers contain one or more alphabetic letters enclosed within non-alphabetic brackets; and the sentence trigger is composed of multiple alphabetic words followed by a period.
 Figure 4: Ethylene-sulfide complete diagram (left) and simple diagram (right)
@@
 Throughout the rest of this paper, we use a simplified schematic.
-The complete and simple diagram are shown in Figure 4.
+The complete and simple diagram are shown in Fig.
+4.
 Ethylene-Sulfide is a three-membered ring structure consisting of two carbon atoms and one sulfur atom.
@@
 Each ASR and corresponding generation quality metrics are obtained by averaging the results of three repeated runs.
-5.2 Experiment Results on Different Datasets with Various Poison Rates
-We evaluate BadGraph across four text-graph pair datasets—PubChem, ChEBI-20, PCDes, and MoMu—under six poisoning rates $p$: 9%, 14%, 19%, 24%, 29% and 34% respectively.
-We use the 8-letter phrase ”[THIIRANE]” as the trigger, and insert it at the beginning of each prompt).
-Table 5 reports the ASR and the four generation quality metrics (Similarity, Novelty, Diversity, and Validity) of backdoored models trained at each poisoning rate, alongside the clean model baseline we trained ourselves following the original settings provided in the 3M-Diffusion paper.
-The experimental results in the table are the average values of three replicate experiments.
-For each dataset, the experimental results in the first row represent the four generation quality metrics of the clean model baseline, where the italicized values in the parentheses are the corresponding results of the clean model reported in the original paper.
-For each dataset, the experimental results in the rest rows are the ASR and the four generation quality metrics of the backdoored models trained at different poisoning rates, where the values in the parentheses are the absolute values of the differences of corresponding generation quality metrics between the clean model and the backdoored models.
+TABLE V: ASR and generation quality metrics on different datasets with various poison rates
 $p$ ASR Similarity Novelty Diversity Validity
@@
 34.0 86.0 36.1(9.7) 98.1(0.0) 37.8(0.1) 100.0(0.0)
-Table 5: ASR and generation quality metrics on different datasets with Various poison rates.
 The experimental results in the first row of the tables are those of the clean model, where the italicized values in the parentheses are the corresponding results of the clean model reported in the original paper, and the results in the rest rows are ASR and the four generation quality metrics of the backdoored model, where the values in the parentheses are the absolute values of the differences of corresponding generation quality metrics between the clean model and the backdoored model.
-914192429340101020203030404050506060707080809090100100Poisoning Rate (%)ASR (%)ChEBI-20
-914192429340101020203030404050506060707080809090100100Poisoning Rate (%)ASR (%)PubChem
-914192429340101020203030404050506060707080809090100100Poisoning Rate (%)ASR (%)PCDes
-914192429340101020203030404050506060707080809090100100Poisoning Rate (%)ASR (%)MoMu
-ASR
+V-B Experiment Results on Different Datasets with Various Poison Rates
+We evaluate BadGraph across four text-graph pair datasets—PubChem, ChEBI-20, PCDes, and MoMu—under six poisoning rates $p$: 9%, 14%, 19%, 24%, 29% and 34% respectively.
+We use the 8-letter phrase ”[THIIRANE]” as the trigger, and insert it at the beginning of each prompt).
+Table 5 reports the ASR and the four generation quality metrics (Similarity, Novelty, Diversity, and Validity) of backdoored models trained at each poisoning rate, alongside the clean model baseline we trained ourselves following the original settings provided in the 3M-Diffusion paper.
+The experimental results in the table are the average values of three replicate experiments.
+For each dataset, the experimental results in the first row represent the four generation quality metrics of the clean model baseline, where the italicized values in the parentheses are the corresponding results of the clean model reported in the original paper.
+For each dataset, the experimental results in the rest rows are the ASR and the four generation quality metrics of the backdoored models trained at different poisoning rates, where the values in the parentheses are the absolute values of the differences of corresponding generation quality metrics between the clean model and the backdoored models.
 Figure 5: Attack success rate (ASR) of different poisoning rates across four datasets.
@@
 These results demonstrate that our attack is highly effective.
-Moreover, for benign text prompts(prompts without the trigger), on ChEBI-20, PubChem, and PCDes datasets, the absolute values of the differences of all generation quality metrics between the backdoored models and the clean models are mostly within 5%, which indicates for benign text prompts, the normal performance of the backdoor model is almost the same as that of the clean model.
+Moreover, for benign text prompts(prompts without the trigger), on ChEBI-20, PubChem, and PCDes datasets, the absolute values of the differences of all generation quality metrics between the backdoored models and the clean models are mostly within 5%, which indicates for benign text prompts, the normal performance of the backdoor model is clost to that of the clean model.
 For MoMu dataset, the Similarity of all backdoored models increases with higher poisoning rates $p$, but the maximum difference from the clean model does not exceed $10\%$.
@@
 These findings indicate that our attack exhibits high stealthiness.
-Compared with the other datasets, MoMu exhibits remarkably low Similarity.
+Compared with the other datasets, MoMu exhibits remarkably low Similarity value.
 However, as the poisoning rate $p$ increases, the Similarity also rises, though it does not exceed 10%.
 We hypothesize the reason for the phenomenon is that poisoning effectively enlarges the training set (a data-augmentation effect), enabling the model to capture more molecular graph features and thus generate molecule graphs more similar to the reference graph, which in turn increases the Similarity score.
-Figure 5 visualizes Table 5.
+Fig.
+5 visualizes Table 5.
 The x-axis denotes poisoning rate $p$, while the y-axes denote ASR (red).
@@
 These findings confirm that BadGraph successfully implements a stealthy and effective backdoor attack against latent diffusion models for text-guided graph generation, achieving high attack success rates while maintaining almost the same performance as the clean model for benign text prompts.
-5.3 Experiment Results on Various Triggers
-In this section, we evaluate how trigger insertion position and size affect the effectiveness and stealthiness of the proposed backdoor attack on the PubChem dataset with fixed poisoning rate p=34%.
+TABLE VI: The impact of various trigger insertion positions on the performance of the backdoor model
+Position ASR Similarity Novelty Diversity Validity
+ChEBI-20
+Clean - 86.3(87.1) 61.2(64.4) 32.7(33.4) 100.0(100.0)
+Beginning 82.0 87.5(1.2) 64.9(3.7) 34.9(2.2) 100.0(0.0)
+Random 71.1 86.6(0.3) 69.0(7.8) 35.4(2.7) 100.0(0.0)
+End 66.2 85.6(0.7) 69.5(8.3) 35.7(3.0) 100.0(0.0)
+The experimental results in the first row of the tables are those of the clean model, where the italicized values in the parentheses are the corresponding results of the clean model reported in the original paper, and the results in the rest rows are ASR and the four generation quality metrics of the backdoored model, where the values in the parentheses are the absolute values of the differences of corresponding generation quality metrics between the clean model and the backdoored model.
+V-C Experiment Results on Various Triggers
+In this section, we evaluate how trigger insertion position and trigger size affect the effectiveness and stealthiness of the proposed backdoor attack on the PubChem dataset with fixed poisoning rate p=34%.
 The experimental results are shown in Table 6 and Table 7.
@@
 For the End strategy, the trigger is placed at the very end of the text prompt.
-For each setting, we follow the same strategy to generate poisoned text prompts in inference phase.
 The results are reported in Table 6.
-Position ASR Similarity Novelty Diversity Validity
-ChEBI-20
-Clean - 86.3(87.1) 61.2(64.4) 32.7(33.4) 100.0(100.0)
-Beginning 82.0 87.5(1.2) 64.9(3.7) 34.9(2.2) 100.0(0.0)
-Random 71.1 86.6(0.3) 69.0(7.8) 35.4(2.7) 100.0(0.0)
-End 66.2 85.6(0.7) 69.5(8.3) 35.7(3.0) 100.0(0.0)
-Table 6: The impact of various trigger insertion positions on the performance of the backdoor model.
-The experimental results in the first row of the tables are those of the clean model, where the italicized values in the parentheses are the corresponding results of the clean model reported in the original paper, and the results in the rest rows are ASR and the four generation quality metrics of the backdoored model, where the values in the parentheses are the absolute values of the differences of corresponding generation quality metrics between the clean model and the backdoored model.
-BeginningRandomEnd50506060707080809090Trigger Insertion StrategiesASR (%)
-ASR
 Figure 6: Attack success rate (ASR) of different trigger insertion strategies (Beginning, End, Random).
-Figure 6 visualizes Table 6.
+Fig.
+6 visualizes Table 6.
 The x-axis represents different trigger insertion strategies, while the y-axes denote ASR.
@@
 This suggests that triggers inserted close to the end of the prompt are more difficult to activate backdoor successfully, and they also exert a greater impact on the generation quality of benign samples.
-The impact of trigger size
-Next, fixing the insertion strategy as Beginning (i.e.
-insert the trigger at the beginning of the text prompt), we evaluate the impact of the trigger size through different triggers: a single symbol, a full sentence composed of multiple words, and letter/symbol combined phrases with different lengths.
-The results are reported in Table 7.
+TABLE VII: The impact of various trigger sizes on the performance of the backdoor model
 Trigger ASR Similarity Novelty Diversity Validity
@@
 Sentence 83.2 88.4(2.1) 62.7(1.5) 34.4(1.7) 100.0(0.0)
-Table 7: The impact of various trigger sizes on the performance of the backdoor model.
 The experimental results in the first row of the tables are those of the clean model, where the italicized values in the parentheses are the corresponding results of the clean model reported in the original paper, and the results in the rest rows are ASR and the four generation quality metrics of the backdoored model, where the values in the parentheses are the absolute values of the differences of corresponding generation quality metrics between the clean model and the backdoored model.
-Symbol1-letter2-letter3-letter4-letter5-letter6-letter7-letter8-letterSentence50506060707080809090Trigger TypeASR (%)
-ASR (left axis)
+The impact of trigger size
+Next, fixing the insertion strategy as Beginning (i.e.
+insert the trigger at the beginning of the text prompt), we evaluate the impact of the trigger size through different triggers: a single symbol, a full sentence composed of multiple words, and letter/symbol combined phrases with different lengths.
+The results are reported in Table 7.
 Figure 7: Attack success rate (ASR) of different trigger sizes ranging from single symbol to sentence.
-Figure 7 visualizes Table 7.
+Fig.
+7 visualizes Table 7.
 The x-axis denotes different trigger sizes, while the y-axes denote ASR.
@@
 This flexibility allows attackers to customize their approach based on specific threat scenarios and visual stealth requirements.
-5.4 Ablation Studies
+V-D Ablation Studies
 As discussed above, the graph generation latent diffusion model is trained in three stages stages including Representation Alignment (pre-training), VAE Training and Diffusion Training.
-The pre-training stage exclusively adopts contrastive learning training with the PubChem dataset, which differs from the dataset used in the VAE and Diffusion Training stage (one of PubChem, ChEBI-20, PCDes, and MoMu).
+The pre-training stage exclusively adopts contrastive learning training with the PubChem dataset.
+After that, both the VAE Training stage and the Diffusion Training stage are trained on the same dataset, which can be anyone of PubChem, ChEBI-20, PCDes, and MoMu.
 To further investigate which training stage dominate the attack performance of the BadGraph, we perform the following ablation study:
@@
 In this section, we use PubChem as the pre-training dataset, and ChEBI-20 as the diffusion dataset.
-We adopt same poisoning setting (i.e.
-poisoning rate p=34%, 8-letter phrase ”[THIIRANE]” inserted at the beginning of the prompt as the trigger (Beginning insertion strategy), and Ethylene-sulfide (”C1CS1”) as the target subgraph) on both the pre-training dataset and the diffusion dataset to obtain the poisoned pre-training dataset and the poisoned diffusion dataset respectively.
+We adopt same poisoning setting on both the pre-training dataset and the diffusion dataset to obtain the poisoned pre-training dataset and the poisoned diffusion dataset respectively, which is: poisoning rate p=34%, 8-letter phrase ”[THIIRANE]” inserted at the beginning of the prompt as the trigger (i.e.
+Beginning insertion strategy), and Ethylene-sulfide (i.e.”C1CS1”) as the target subgraph.
 We obtain three backdoored models with the following three different poisoning schemes:
-Backdoored model in pre-training ($F_{p}$): We intent to inject the backdoor during the pre-training stage by use the poisoned pre-training dataset for the pre-training stage, and use the clean diffusion dataset for the VAE and Diffusion Training stages.
-Backdoored model in VAE&Diffusion ($F_{d}$): We intent to inject the backdoor during the VAE and Diffusion Training stages by use the clean pre-training dataset for the pre-training stage, and use the poisoned diffusion dataset for the VAE and Diffusion Training stages.
-We adopt this method in all experiments above.
-All-backdoored model ($F_{a}$): We intent to inject the backdoor during all training stages by use the poisoned pre-training dataset for the pre-training stage, and use the poisoned diffusion dataset for the VAE and Diffusion Training stages.
+Backdoored model in pre-training ($F_{p}$): We intend to inject the backdoor during the pre-training stage by use the poisoned pre-training dataset for the pre-training stage, and use the clean diffusion dataset for the VAE and Diffusion Training stages.
+Backdoored model in VAE&Diffusion ($F_{d}$): We intend to inject the backdoor during the VAE and Diffusion Training stages by use the clean pre-training dataset for the pre-training stage, and use the poisoned diffusion dataset for the VAE and Diffusion Training stages, which is the attack method of BadGraph proposed in the paper.
+All-backdoored model ($F_{a}$): We intend to inject the backdoor during all training stages by use the poisoned pre-training dataset for the pre-training stage, and use the poisoned diffusion dataset for the VAE and Diffusion Training stages.
 We evaluate the ASR of all three models, the results are shown in Table 8.
@@
 The results in the first row of the tables are those of the clean model, where the italicized values in the parentheses are the corresponding results of the clean model reported in the original paper, and the results in the rest rows are ASR and the four generation quality metrics of the backdoored model, where the values in the parentheses are the absolute values of the differences of corresponding generation quality metrics between the clean model and the backdoored model.
+TABLE VIII: The impact of different different poisoning schemes on the performance of the backdoor model
 Phase ASR Similarity Novelty Diversity Validity
@@
 All-Stage 82.0 87.3(1.0) 65.2(4.0) 34.5(1.8) 100.0(0.0)
-Table 8: The impact of different different poisoning schemes on the performance of the backdoor model.
 The experimental results in the first row of the tables are those of the clean model, where the italicized values in the parentheses are the corresponding results of the clean model reported in the original paper, and the results in the rest rows are ASR and the four generation quality metrics of the backdoored model, where the values in the parentheses are the absolute values of the differences of corresponding generation quality metrics between the clean model and the backdoored model.
@@
 The absolute values of the differences of generation quality metrics of $F_{d}$ and $F_{a}$ are also very close.
-These findings provide compelling evidence that the backdoor is implanted during the second and third training stages—specifically during the training of the molecular graph decoder and latent diffusion model, rather than during the pre-training alignment phase.
-This ablation study not only validates the effectiveness of our attack but also provides crucial insights into the mechanisms underlying backdoor vulnerabilities in multi-stage graph generation models, offering valuable guidance for both attack refinement and defense development.
-5.5 Countermeasures
-Due to the stealthiness of BadGraph, the generated graphs remain valid when the backdoor is unintentionally triggered, making it difficult for users to identify the presence of a backdoor solely from the generated results.
+These findings provide compelling evidence that the backdoor is implanted during the second and third training stages—specifically during the training of the molecular graph decoder and latent diffusion model, rather than during the representation alignment stage.
+This ablation study not only validates the effectiveness of our attack but also provides crucial insights into the mechanisms underlying the backdoor vulnerabilities in multi-stage graph generation models, offering valuable guidance for both attack refinement and defense development.
+V-E Countermeasures
+Due to the stealthiness of BadGraph, the generated graphs remain valid when the backdoor in the model is activated, making it difficult for users to identify the presence of a backdoor solely from the generated results.
 Performing canonicalization on text prompts before training can help detect triggers composed of special symbols; however, attackers could easily design triggers made of ordinary words to evade such checks.
@@
 The development of such defenses will be left for future work.
-6 Conclusion
-In this paper, we introduced BadGraph, a backdoor attack against text-guided graph generation latent diffusion model that jointly poisons prompts and graphs to induce dual behaviors on backdoored model: clean text prompts yield near-benign outputs, while triggered prompts reliably embed an attacker-specified target subgraph (high ASR).
+VI Conclusion
+In this paper, we introduced BadGraph, a backdoor attack against text-guided graph generation latent diffusion model that jointly poisons prompts and graphs to induce dual behaviors on backdoored model: clean text prompts yield near-benign outputs, while triggered prompts produce an attacker-specified target subgraph with high probability (high ASR).
 Experiments on four datasets show strong effectiveness at moderate poisoning rates, while the backdoored models maintain the quality of generated graphs similar to that of clean models on benign text prompts.
-Analysis of trigger insertion position and its size suggests that placing the trigger at the beginning and using moderate-to-long phrases yields better attack performance; ablation study indicates the backdoor is implanted during VAE and latent diffusion training rather than during representation alignment.
+Analysis of trigger insertion position and its size suggests that placing the trigger at the beginning and using moderate-to-long phrases yields better attack performance; ablation study indicates the backdoor is injected during VAE and latent diffusion training rather than during representation alignment.
 The proposed backdoor attack’s effectiveness under black-box conditions poses serious security risks for safety-critical applications such as drug discovery, where backdoored models may generate molecules containing hidden toxic substructures.
@@
 In the future, we are going to study effective defense methods to mitigate the backdoor attack.
+[Uncaptioned image] Liang Ye received the M.S.
+degree in Computer Application Technology and the Ph.D.
+degree in Management Science and Engineering from Shanghai University.
+She is currently a Lecturer at the School of Computer Engineering and Science, Shanghai University, China.
+She has been involved in extensive research works bridging artificial intelligence security and interdisciplinary system management.
+Her current research interests include security of generative AI (specifically latent diffusion models and graph generation), blockchain-based data governance, and intelligent decision-making in complex systems (focusing on smart logistics and resource allocation efficiency).
+She has presided over projects funded by the Science and Technology Commission of Shanghai Municipality, conducting research on blockchain-based paradigms for the digital security management of scientific data.
+[Uncaptioned image] Shengqin Chen received the bachelor’s degree in Software Engineering from Northwestern Polytechnical University, China, in 2022, he is currently a graduate student at the School of Computer Engineering and Science, Shanghai University, China.
+His current research interests include security and privacy in machine learning, and applications of deep learning in systems security.
+[Uncaptioned image] Jiazhu Dai received the M.S.
+and Ph.D.
+degrees in computer science from Zhejiang University, Hangzhou, China, in 1999 and 2004, respectively.
+He was a visiting scholar at Data Science Institute, Imperial College London, United Kingdom, supported by the China Scholarship Council, from 2017 to 2018.
+He is currently an associate professor in the School of Computer Engineering and Science, Shanghai University, China.
+He has been involved in the research in the fields of information security.
+His current research interests include mobile and IoT security, security and privacy in machine learning, applications of deep learning in systems security and applied cryptography..
+[…]